### v1

# Generative Autoregressive Transformers for Model-Agnostic Federated MRI Reconstruction

Abstract
Although learning-based models hold great promise for MRI reconstruction, single-site models built on limited local datasets often suffer from poor generalization. This challenge has spurred interest in collaborative model training on multi-site datasets via federated learning (FL)– a privacy-preserving framework that aggregates model updates instead of sharing imaging data. Conventional FL builds a global model by aggregating locally trained model weights, inherently constraining all sites to a homogeneous model architecture. This rigid homogeneity requirement forces sites to forgo architectures tailored to their compute infrastructure and application-specific demands. Consequently, existing FL methods for MRI reconstruction fail to support model-heterogeneous settings, where individual sites are allowed to use distinct architectures. To overcome this fundamental limitation, here we introduce FedGAT, a novel model-agnostic FL technique based on generative autoregressive transformers. FedGAT decentralizes the training of a global generative prior that captures the distribution of multi-site MR images. For enhanced fidelity, we propose a novel site-prompted GAT prior that controllably synthesizes MR images from desired sites via autoregressive prediction across spatial scales. Each site then trains its site-specific reconstruction model–using its preferred architecture–on a hybrid dataset comprising the local MRI dataset and GAT-generated synthetic MRI datasets for other sites. Comprehensive experiments on multi-institutional datasets demonstrate that FedGAT supports flexible collaborations while enjoying superior within-site and across-site reconstruction performance compared to state-of-the-art FL baselines.
{IEEEkeywords}
MRI, reconstruction, federated learning, model agnostic, generative, autoregressive, transformer.
[…]
IEEEexample:BSTcontrol
1 Introduction
[…]
Magnetic Resonance Imaging (MRI) is an indispensable diagnostic modality due to its exceptional soft-tissue contrast. Yet, characteristically long scan times in MRI exams can disrupt clinical workflows and reduce patient comfort [1]. A mainstream approach to improve scan efficiency relies on reconstruction of MR images from undersampled acquisitions [2, 3]. Learning-based reconstruction models have shown particular promise in recovering high-quality images from undersampled k-space data [4, 5, 6, 7, 8, 9, 10, 11, 12]. Despite their potential, models trained on local datasets available at individual sites often struggle to represent rare tissue features and fail to generalize effectively across sites [13, 14, 15, 16, 17, 18, 19, 20, 21]. As such, the development of robust models remains hindered by the limited availability of diverse training datasets in solitary institutions due to economic costs and privacy concerns [22, 23, 24]. This limitation underscores the pressing need for collaborative approaches to build models on multi-site MRI datasets while preserving patient privacy [25, 26].
Federated learning (FL) is an emergent paradigm for collaborative model training in privacy-sensitive domains such as medical imaging [27, 28]. By exchanging model weights as opposed to raw imaging data, FL enables knowledge transfer among sites while alleviating potential privacy risks [29]. Recognizing the transformative potential of FL, several recent studies on MRI reconstruction have proposed FL methods for building models on multi-site datasets [30, 31, 32, 33, 34, 35]. A common theme among these studies has been their adherence to the conventional FL framework based on a server-client topology [29]. In this framework, a globally shared model is selected via agreement across sites to serve as the foundation for both cross-site knowledge transfer and eventual execution of the reconstruction task. Over multiple communication rounds, local model copies trained on local datasets available at individual sites are aggregated on the server to update the global model [36]. The global model is eventually expected to possess knowledge on the distribution of multi-site MRI datasets and thereby improve generalization over single-site models trained exclusively on local datasets [30].
The conventional FL framework promises enhanced generalization in MRI reconstruction, while imposing a strict requirement for all participating sites to adopt a homogeneous model architecture as the basis of model aggregation [36]. However, imaging sites often differ substantially in terms of their computational resources and the complexity of their reconstruction tasks as influenced by variations in distribution of local datasets. These differences frequently lead sites to favor distinct model architectures, ranging from convolutional networks that can be preferred to attain high local precision under relatively low compute budgets [30, 31], transformers that can be preferred to boost sensitivity to long-range contextual features when high compute budgets are available [37, 38], to physics-driven unrolled networks that can be preferred to boost model reliability when training sets are relatively scarce albeit moderate-to-high compute budgets are available [34, 33]. Consequently, the architectural constraints in conventional FL refrain sites from leveraging models optimized for their specific requirements. This apparent lack of architectural freedom has important practical implications including discouraging resource-constrained sites from participating in collaborations, compromising performance at sites with more advanced resources due to use of relatively simple models, and ultimately stifling innovations in MRI reconstruction models. Therefore, enabling model-agnostic FL—where sites can collaborate without adhering to a homogeneous model architecture—remains a critical open challenge [39].
In this study, we introduce a novel FL technique based on generative autoregressive transformers (FedGAT), which to our knowledge is the first model-agnostic FL approach in the literature for MRI reconstruction. Unlike conventional FL methods, FedGAT supports model-heterogeneous settings by decoupling the process of cross-site knowledge transfer on data distributions, from the process of building reconstruction models (Fig. 1). It mediates knowledge transfer via a global generative prior designed to synthesize multi-site MR images. To avoid the limitations of existing adversarial priors (i.e., training instability and poor quality) and diffusion priors (i.e., suboptimal convergence and heavy compute burden), here we propose a novel GAT prior that autoregressively synthesizes MR images by predicting feature maps across spatial scales, guided by a site prompt to preserve site-specific characteristics. Afterwards, site-specific reconstruction models are locally trained on a hybrid dataset, combining local MRI datasets available at each site with synthetic datasets from remaining sites generated by the GAT prior. Comprehensive demonstrations on multi-institutional MRI datasets indicate that FedGAT functions effectively in model-heterogeneous settings, overcoming a significant challenge in enabling privacy-preserving collaborations across diverse institutions. Code for FedGAT is available at https://github.com/icon-lab/FedGAT.
Contributions
  • To the best of our knowledge, FedGAT is the first model-agnostic federated learning (FL) method for MRI reconstruction, enabling flexible collaborations among sites with heterogeneous model preferences.
  • To support model heterogeneity, FedGAT decouples decentralized training of a global generative prior on multi-site MR images for cross-site knowledge transfer, from local training of site-specific reconstruction models on a combination of local and synthetic MRI datasets.
  • To improve synthesis fidelity, FedGAT employs a novel prior based on generative autoregressive transformers, formulating image generation as an autoregressive prediction task across spatial scales and enforcing control over the data distribution via a site-prompting mechanism.
  • To strike a balance between generalization and site-specific performance, FedGAT leverages a training procedure where reconstruction models pre-trained on local datasets are fine-tuned on a hybrid dataset including synthetic datasets from other sites.
Figure 1: FedGAT devises a two-tier strategy to collaboratively train heterogeneous models for MRI reconstruction. (a) The first tier conducts decentralized training of a global prior $\text{GAT}_{\theta_{\text{GAT}}}$ that captures the distribution of multi-site MR images. The global prior is built on a novel generative autoregressive transformer that formulates image generation as an autoregressive process across growing spatial scales, and that employs a site prompt sp for controllable image synthesis from desired sites. (b) The second tier conducts local training of site-specific reconstruction models $H^{k}_{{\phi}^{k}}$ ($k$: site index). To achieve an optimal balance between generalization and site-specific performance, training at each site is performed on a hybrid dataset comprising the local dataset and synthetic datasets from other sites generated via the GAT prior.
2 Related Work
2.1 Learning-Based MRI Reconstruction
Learning-based models promise improved performance and reliability over traditional methods in MRI reconstruction [5]. Yet, realizing this promise requires copious amounts of training data including diverse samples, since learning-based models poorly represent rare features and pathologies in their training sets [40]. Unfortunately, curating broad training sets is unfeasible in many applications given the high costs of scanning and subject recruitment [41]. These costs bear the common practice of model training on limited datasets available locally in solitary imaging sites, and such single-site models can inevitably suffer from suboptimal generalization [42]. Numerous strategies have been devised over the years to alleviate the unwanted influences of dataset scarcity, including transfer learning [7], unpaired learning [15, 43], semi-supervised learning [44], and self-supervised learning [45, 46]. However, these canonical strategies rely on centralized model training following aggregation of datasets in a central repository, which incurs significant privacy concerns and maintenance costs [25].
2.2 Federated Learning for MRI Reconstruction
Federated learning (FL) has emerged as a privacy-preserving collaboration framework that distributes the costs of model training across healthcare institutions [25]. Abandoning explicit sharing of sensitive imaging data, FL promotes cross-site knowledge transfer via exchange of model weights instead. In conventional FL, this exchange is attained by first training copies of a global model on the local datasets available at individual sites, and then aggregating the local copies into the shared global model on a server [27]. Several recent MRI studies have successfully adopted this conventional FL framework to boost generalization performance of reconstruction models [30, 31, 32]. Further improvements in site-specific performance have also been sought by employing personalized FL strategies to cope with differences in the distribution of MRI datasets across sites, such as partial model aggregation [32], test-time adaptation [31], and feature map normalization [42].
Despite their apparent benefits, previous FL methods for MRI reconstruction constrain all participating sites to adopt a homogeneous model architecture, which serves as the foundation for aggregating model weights [36]. Note, however, that imaging sites can have substantial differences in computational resources or in difficulty of reconstruction tasks dependent on the interaction between distribution of local MRI datasets and desired acceleration rates. In turn, these differences can drive individual sites to prefer distinct models for MRI reconstruction, as evident from the literature where model preferences range broadly from convolutional [30, 41, 32] and transformer [38, 47, 37] backbones to physics-driven unrolled architectures [14, 34, 33, 48]. As such, the rigid model-homogeneity requirement of conventional FL severely limits the flexibility of individual sites, forcing them to forgo architectures tailored to their specific needs. During multi-institutional collaborations, this restriction can hinder participation of sites with limited resources or compromise performance in sites with copious resources.
To enable collaborative training of heterogeneous models across sites, here we introduce FedGAT as the first model-agnostic FL method for MRI reconstruction. Unlike conventional FL that transfers knowledge by communicating weights of the reconstruction model, FedGAT leverages a unique approach that decouples the process of knowledge transfer across sites from the process of building site-specific reconstruction models, and that mediates knowledge transfer via a global generative prior for multi-site MR images. Note that previous FL methods in machine learning that build generative priors have commonly proposed adversarial priors that can suffer from poor training stability and image quality [49, 31], and diffusion priors that can suffer from suboptimal convergence that can lead to residual image noise and prolonged run times [50]. In contrast, here we introduce a novel prior based on generative autoregressive transformers that efficiently synthesizes MR images via autoregressive prediction of feature maps across growing spatial scales, under guidance from a site prompt to preserve site-specific characteristics in MR images. At each site, a site-specific reconstruction model is then locally trained on a hybrid dataset comprising both the local dataset and synthetic datasets from other sites generated via the global prior. These technical advances enable FedGAT to operate seamlessly in model-heterogeneous settings, overcoming a critical barrier towards privacy-preserving collaborations across diverse institutions.
3 Theory
3.1 Conventional FL for MRI Reconstruction
In MRI, the image domain that depicts the spatial distribution of tissues and the measurement domain where k-space data are acquired are linked through an imaging operator $\mathcal{A}=\mathcal{M}\mathcal{F}\mathcal{C}$ ($\mathcal{M}$: sampling pattern, $\mathcal{F}$: Fourier transform, $\mathcal{C}$: coil sensitivities):
$\mathcal{A}\mathbf{x}=\mathbf{y},$ (1)
where $\mathbf{x}$ is the MR image and $\mathbf{y}$ are acquired k-space data. For undersampled acquisitions, adequate solution of Eq. 1 requires incorporation of a regularizing constraint $\mathcal{R}(\cdot)$ [1]:
$\widehat{\mathbf{x}}=\underset{\mathbf{x}}{\operatorname{argmin}}\|\mathbf{y}-\mathcal{A}\mathbf{x}\|_{2}^{2}+\mathcal{R}(\mathbf{x}).$ (2)
Over the recent years, learning-based methods have gained prominence that operationalize the solution of Eq. 2 as projection through a network model, typically operating in image domain as $\widehat{\mathbf{x}}=H_{\phi}(\mathcal{F}^{-1}\mathbf{y},\mathcal{A})$, which receives a zero-filled Fourier reconstruction of undersampled data as input along with the imaging operator [5].
In the conventional FL framework, an MRI reconstruction model is collaboratively trained over $N_{c}$ rounds of communication between a server and participating sites [27, 28]. The server maintains a global reconstruction model $H_{\phi}$ with weights $\phi$, and sites retain local copies of this common architecture $\{H_{\phi^{k}}\}_{k=1}^{K}$ where $k$ is site index, $K$ is the number of sites. At the start of the $c$th communication round, all sites initialize their local copies with the weights of the global model broadcast by the server, $\phi^{k}\leftarrow\phi(c-1)$. At the $k$th site, the initialized local copy $H_{\phi^{k}}$ is then trained on the locally available dataset $\mathcal{D}_{\text{loc}}^{k}$ to minimize a reconstruction loss:
$\mathcal{L}^{k}_{\text{rec}}(\mathcal{D}_{\text{loc}}^{k})=\mathbb{E}_{(\mathbf{x}^{k},\mathbf{y}^{k})\sim\mathcal{D}_{\text{loc}}^{k}}\left[\|\mathbf{x}^{k}-H_{\phi^{k}}(\mathcal{F}^{-1}\mathbf{y}^{k},\mathcal{A}^{k})\|_{2}^{2}\right],$ (3)
where $\mathbb{E}$ denotes expectation over $\mathcal{D}_{\text{loc}}^{k}$, $\mathbf{x}^{k}$ is the reference image derived from a fully-sampled acquisition, $\mathbf{y}^{k}$ is the corresponding undersampled acquisition, and $\mathcal{A}^{k}$ is the imaging operator. At the end of the given communication round, locally-trained model copies are aggregated on the server via weight averaging to derive an updated global model [36]:
$\phi(c)=\sum_{k=1}^{K}\alpha_{k}\phi^{k},$ (4)
where $\alpha_{k}=\frac{N^{k}}{N^{1}+...+N^{K}}$ denotes the relative contribution of the model copy from site $k$ ($N^{k}$: the number of training samples at site $k$).
After a total of $N_{c}$ communication rounds, the resultant global model $H_{\phi^{*}}$ with $\phi^{*}:=\phi(N_{c})$ is employed at each site to reconstruct MR images. As such, reconstruction for a test subject at site $k$ is formulated as:
$\widehat{\mathbf{x}}^{k}=H_{\phi^{*}}(\mathcal{F}^{-1}\mathbf{y}^{k},\mathcal{A}^{k}),$ (5)
where $\widehat{\mathbf{x}}^{k}$ is the reconstructed image, and $\mathbf{y}^{k}$ are the respective undersampled k-space data.
3.2 Federated Generative Autoregressive Transformers
FedGAT is a model-agnostic FL technique that enables collaborative training of heterogeneous model architectures across sites by leveraging a decoupled two-tier strategy. During the first tier, using a server-client topology, a site-prompted global prior based on generative autoregressive transformers (GAT) is decentrally trained so as to capture the distribution of multi-site MR images (Fig. 1a). The FL server sends the trained global prior to individual sites at the end of the first tier. During the second tier, site-specific reconstruction models are locally trained with the aid of this global GAT prior (Fig. 1b). At each site, a reconstruction model of preferred architecture is pre-trained on the local dataset, and later fine-tuned on a hybrid dataset containing both the local dataset and synthetic datasets from other sites generated via the prior. The architecture of the site-prompted GAT prior, and image synthesis and two-tier training procedures for FedGAT are detailed in subsections below.
3.2.1 Architecture of the Site-Prompted GAT Prior
FedGAT uses a global generative prior to capture the distribution of multi-site MR images, such that synthetic MR images from individual sites can later be generated (Fig. 2). Given the recent success of autoregressive models in computer vision tasks on natural images [51] and the unparalleled contextual sensitivity of transformers in MR image formation tasks [47, 52], here we propose a novel GAT prior equipped with transformers to autoregressively generate multi-site MR images. Unlike traditional autoregressive models devised to predict the next image token given earlier tokens (i.e., image pixels or small image patches) [53], the autoregressive process in GAT predicts feature maps at higher spatial scales given maps at lower scales, as inspired by a recent study on natural images [54]. Yet, differently from [54], GAT is designed for generating multi-site MRI data by introducing a novel site-prompting mechanism that controls the sites from which MR images are synthesized.
Since transformer architectures process images as a sequence of tokens under quadratic complexity with respect to sequence length, using transformers on pixel-level tokens is computationally challenging, whereas using them on larger patch-level tokens can degrade spatial precision [55]. To mitigate computational burden without compromising spatial precision, here we leverage autoregressive transformers to instead synthesize feature maps in a compact latent space, and map between the image domain and the latent space via a variational autoencoder [56], drawing inspiration from recent efforts in efficient generative modeling [57]. Thus, the GAT prior uses a compound architecture with VAE encoder-decoder and autoregressive transformer modules.
VAE Encoder Module: Receiving a coil-combined complex MR image $\mathbf{x}\in\mathbb{R}^{H\times W\times 2}$ with real and imaginary components stored as separate channels in the third tensor dimension, the VAE encoder aims to map its input onto discrete token maps $\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{S}$ across $S$ spatial scales. To do this, the encoder first extracts a continuous latent representation $\mathbf{z}\in\mathbb{R}^{h_{S}\times w_{S}\times c}$ ($h_{S}$, $w_{S}$ denoting spatial dimensions at the $S$th scale–the highest scale) via a set of residual convolutional blocks: $\mathbf{z}=\mathrm{ResConv}(\mathbf{x})$ [58]. Afterwards, the discrete token maps at the $s$th scale are extracted by spatial downsampling to attain dimensions of $h_{s}\times w_{s}$, followed by a quantization procedure based on a learnable codebook $\mathbf{B}\in\mathbb{R}^{V\times c}$ with vocabulary size $V$. Note that $\mathbf{B}$ characterizes a discrete latent space comprising $V$ categories.
To minimize redundancy and information losses across scales, here we adopt a hierarchical procedure for token map extraction that progresses from the lowest ($s$=\,=\,=$1$) to the highest ($s$=\,=\,=$S$) spatial scale [59]. For this purpose, a residual continuous representation $\mathbf{r}_{s}\in\mathbb{R}^{h_{S}\times w_{S}\times c}$ is maintained, initialized as $\mathbf{r}_{1}=\mathbf{z}$ at $s$=\,=\,=$1$. At the $s$th scale, the residual continuous representation is used to derive $\mathbf{f}_{s}$ as follows:
$\mathbf{f}_{s}=\underset{v\in\{1,\dots,V\}}{\mathrm{argmin}}\,\lVert\mathbf{B}(v\,,\,:)-\text{Down}_{s}(\mathbf{r}_{s})\rVert^{2}_{2},$ (6)
where $\mathbf{f}_{s}\in[V]^{h_{s}\times w_{s}}$, $\text{Down}_{s}$ denotes spatial downsampling to the $s$th scale via interpolation, and quantization is attained by identifying the closest vector in the codebook according to Euclidean distance. Following [56], the codebook vectors in $\mathbf{B}$ are randomly initialized from a uniform distribution during training, which has been shown to yield a uniform prior on the discrete token maps $\mathbf{f}_{s}$ when accompanied by the quantization objective in Eq. 6. Afterwards, codebook vectors corresponding to $\mathbf{f}_{s}$ are retrieved and upsampled via interpolation to the $S$th scale, and used to update the residual representation:
$\mathbf{r}_{s+1}=\mathbf{r}_{s}-\text{Conv}(\text{Up}_{S}(\text{Lookup}(\mathbf{B},\mathbf{f}_{s}))),$ (7)
where Lookup is the retrieval function for codebook vectors given discrete token maps, $\text{Up}_{S}$ denotes spatial upsampling to the $S$th scale via interpolation, and Conv denotes a convolutional layer. Note that, at the encoder output, the discrete token maps derived at individual spatial scales are pooled into an aggregate multi-scale token map:
$\{\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{S}\}=\mathrm{Enc}(\mathbf{x}),$ (8)
which contains a set of foundational features of MR images.
Figure 2: Architecture of the proposed site-prompted GAT prior. (a) The GAT prior embodies a variational autoencoder (VAE), whose encoder module maps an input MR image onto a set of discrete token maps $\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{S}$ across $S$ spatial scales, and whose decoder module recovers the input MR image back from the derived token maps. (b) The GAT prior also embodies a transformer module that builds an autoregressive prior on the multi-scale discrete token maps captured by the VAE. This module expresses generation of multi-scale discrete token maps as an autoregressive process across spatial scales, wherein a higher-scale token map $\mathbf{f}_{s}$ is predicted given the collection of token maps at earlier scales $\mathbf{f}_{<s}$:=assign\,:=\,:=$\{\mathbf{f}_{1},...,\mathbf{f}_{s-1}\}$. To preserve site-specific attributes in synthetic MR images, a site-token $\mathbf{st}$ derived from a site prompt is included in the autoregressive transformer module.
VAE Decoder Module: The VAE decoder aims to map discrete multi-scale token maps $\{\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{S}\}$ back on the corresponding MR image $\mathbf{x}$ from which they were derived. Paralleling the extraction procedure in the VAE encoder, the decoder employs a hierarchical procedure to recover the image progressively from low-to-high spatial scales [59]. In particular, a residual continuous representation is initialized as $\hat{\mathbf{r}}_{0}=0\in\mathbb{R}^{h_{S}\times w_{S}\times c}$. At the $s$th scale, codebook vectors corresponding to the discrete token map $\mathbf{f}_{s}$ are retrieved and upsampled, and used to update the predictions for the residual representation as follows:
$\hat{\mathbf{r}}_{s}=\hat{\mathbf{r}}_{s-1}+\text{Conv}(\text{Up}_{S}(\text{Lookup}(\mathbf{B},\mathbf{f}_{s}))).$ (9)
Following $S$ scales of processing, the predicted continuous representation $\hat{\mathbf{z}}=\hat{\mathbf{r}}_{S}\in\mathbb{R}^{h_{S}\times w_{S}\times c}$ at the highest scale is used to recover the original image via projection through a set of residual convolutional blocks:
$\hat{\mathbf{x}}=\text{ResConv}(\hat{\mathbf{r}}_{S})=\mathrm{Dec}(\{\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{S}\}),$ (10)
where $\hat{\mathbf{x}}\in\mathbb{R}^{H\times W\times 2}$ denotes a prediction of the original MR image input to the VAE encoder.
Autoregressive Transformer Module: While the VAE encoder extracts discrete token maps residually across spatial scales, it does not explicitly examine the statistical relationships among these multi-scale maps [56]. To obtain a correlate for the joint distribution of $\mathbf{f}_{1},..,\mathbf{f}_{S}$, here we propose to build an autoregressive prior on multi-scale discrete token maps via the autoregressive transformer module. In particular, the transformer module aims to predict the sequence of tokens $\mathbf{f}_{s}$ at the $s$th scale, given as input the sequence of tokens aggregated across lower scales $\mathbf{f}_{<s}$:=assign\,:=\,:=$\{\mathbf{f}_{1},...,\mathbf{f}_{s-1}\}\in[V]^{T_{s-1}}$ ($T_{s-1}=\sum_{o=1}^{s-1}h_{o}w_{o}$ denoting the number of tokens at lower scales), and a learnable site-prompt $\mathbf{sp}(k)\in\mathbb{R}^{1\times d}$ used to initialize a site token, $\mathbf{st}$=\,=\,=$\mathbf{sp}(k)$. To maintain spatial correspondence between the token maps at consecutive scales, the token maps in $\mathbf{f}_{<s}$ are individually embedded after upsampling to the dimensions of the subsequent scale:
$\mathbf{e}_{o}=\mathrm{Lin}(\mathrm{Up}_{o+1}(\mathbf{f}_{o})),\mbox{ for }o\in[1\mbox{ }s-1],$ (11)
where $\mathbf{e}_{o}\in\mathbb{R}^{h_{o+1}w_{o+1}\times d}$ and $\mathbf{e}_{<s}$:=assign\,:=\,:=$\{\mathbf{e}_{1},...,\mathbf{e}_{s-1}\}\in\mathbb{R}^{(T_{s}-1)\times d}$. Afterwards, the site token is pooled with a learnable projection of these embedded token maps, and additively combined with learnable position and scale encodings:
$\mathbf{h}_{0}=\Big{[}\big{[}\mathbf{st};\,\,\mathbf{e}_{<s}\cdot\mathbf{E}_{\text{tok}}\big{]}+\mathbf{E}_{\text{pos}}+\mathbf{E}_{\text{sca}};\,\,\mathbf{m}\Big{]},$ (12)
where $\mathbf{h}_{0}\in\mathbb{R}^{T_{S}\times d}$ is a hidden representation with $T_{S}$ denoting the number of tokens across all scales, $\mathbf{E}_{\text{tok}}\in\mathbb{R}^{(T_{s}-1)\times d}$ is the token projection matrix, $\mathbf{E}_{\text{pos}}\in\mathbb{R}^{T_{s}\times d}$ is the position encoding that captures the relative spatial locations of tokens [60], and $\mathbf{E}_{\text{sca}}\in\mathbb{R}^{T_{s}\times d}$ is the scale encoding that captures the relative spatial scales of tokens across $s$ scales. In Eq. 12, $\mathbf{m}$=\,=\,=$0^{(T_{S}-T_{s})\times d}$ is a zero matrix that masks tokens from spatial scales succeeding the $(s-1)$th scale to ensure causal progression from lower to higher scales.
To predict $\mathbf{f}_{s}$, the hidden representation $\mathbf{h}_{0}$ is projected through $L$ transformer blocks, each containing multi-head self-attention (MHSA) and multi-later perceptron (MLP) layers interleaved with normalization and residual connections [55]. Yet, unlike normalization layers in conventional transformers, here we employ adaptive layer normalization (AdaLN) in order to induce site-specific processing of hidden representations [61]. In particular, we propose to modulate the statistics of $\mathbf{h}$ via learnable functions of the site token $\mathbf{st}$:
$\mathrm{AdaLN}(\mathbf{h},\mathbf{st})=\gamma(\mathbf{st})\cdot\frac{\mathbf{h}-\mu(\mathbf{h})}{\sigma(\mathbf{h})}+\beta(\mathbf{st}),$ (13)
where $\mathbf{h}$ is the hidden representation, $\mu$ and $\sigma$ are the layer mean and standard deviation for $\mathbf{h}$, and $\beta$, $\gamma\,\in\mathbb{R}$ are learnable bias and gain parameters dependent on $\mathbf{st}$. Accordingly, the projection through the $\ell$th transformer block ($\ell\in[1\mbox{ }L]$) can be expressed as follows:
$\mathbf{h}^{\prime}_{\ell}$ $=$ $\mathbf{h}_{\ell-1}+\mathrm{MHSA}(\mathrm{AdaLN}(\mathbf{h}_{\ell-1},\mathbf{st})),$ (14)
$\mathbf{h}_{\ell}$ $=$ $\mathbf{h}^{\prime}_{\ell}+\mathrm{MLP}(\mathrm{AdaLN}(\mathbf{h}^{\prime}_{\ell},\mathbf{st})).$ (15)
In MHSA layers, attention scores between query and key tokens in $\mathbf{h}$ are obtained by computing softmax dot-product similarity [60], and these attention scores are then used to filter the value tokens:
$A_{i,j}=\text{Softmax}\left(\frac{\mathbf{q}_{i}\cdot\mathbf{k}_{j}^{\top}}{\sqrt{d}}\right);\quad\mathrm{SA}(\mathbf{h})=A\cdot\mathbf{v},$ (16)
where $A\in\mathbb{R}^{(1+T_{S})\times(1+T_{S})}$ is the attention score matrix, $i$, $j$ are token indices, $\mathbf{q}$, $\mathbf{k}$, $\mathbf{v}$ denote unit-norm query, key, value tokens obtained as learnable linear projections of tokens in $\mathbf{h}$, and $\mathrm{SA}$ denotes a single self-attention head in MHSA.
After projection through the final transformer block, an adaptive layer norm conditioned on $\mathbf{st}$ followed by a linear layer is employed to extract the logits $\mathbf{logit}\in\mathbb{R}^{(h_{s}w_{s})\times V}$ with improved site specificity, the probability values for these tokens $\mathbf{P}_{s}$ are computed from the logits, and the specific codebook indices maximizing the probability values are drawn to compute the predicted token sequence:
$\mathbf{logit}$ $=$ $\mathrm{Lin}(\mathrm{AdaLN}(\mathbf{h}_{L}(T_{s-1}+1:T_{s},:),\mathbf{st}))$ (17)
$\mathbf{P}_{s}$ $=$ $\mathrm{Softmax}({\mathbf{logit}}),$ (18)
$\hat{\mathbf{f}}_{s}(i)$ $=$ $\underset{v\in\{1,\dots,V\}}{\mathrm{argmax}}\,\mathbf{P}_{s}(i,v),$ (19)
where $\hat{\mathbf{f}}_{s}(i)$ is the prediction for the $i$th token in $\mathbf{f}_{s}$, and $\mathbf{P}_{s}(i,v)$ is the predicted probability of the $v$th codebook vector for $\mathbf{f}_{s}(i)$. For brevity, we will refer to the autoregressive mapping performed by the transformer module as follows:
$\hat{\mathbf{f}}_{s}=\mathrm{Trans}(\mathbf{sp}(k),\,\mathbf{f}_{<s}),\mbox{for }s\in[1\mbox{ }{S}].$ (20)
Note that $\mathbf{f}_{<1}=\{\mathbf{f}_{0}\}\in\varnothing$ does not contain any tokens.
Input: $\mathcal{D_{\text{loc}}}=\{\mathcal{D}_{\text{loc}}^{1},...,\mathcal{D}_{\text{loc}}^{K}\}$: local datasets from $K$ sites.
GAT: global GAT prior with parameters $\theta_{\text{GAT}}$.
$\text{GAT}^{1},...,\text{GAT}^{K}$}: local copies with $\{\theta_{\text{GAT}}^{1},...,\theta_{\text{GAT}}^{K}\}$.
$N_{c}$: number of communication rounds.
$N_{l}$: number of local epochs.
Output: Trained global GAT prior with $\hat{\theta}_{\text{GAT}}$.
1 Initialize global prior with $\theta_{\text{GAT}}(0)$.
2 for $c=1:N_{c}$ do
3       for $k=1:K$ do
4             $\theta_{\text{GAT}}^{k}(0)\leftarrow\theta_{\text{GAT}}(c-1)$triangleright global onto local copy
5             for $e=1:N_{l}$ do
6                   $\theta_{\text{GAT}}^{k}(e)\leftarrow\theta_{\text{GAT}}^{k}(e-1)-\eta\nabla_{\theta_{\text{GAT}}^{k}}\mathcal{L}^{k}_{\text{GAT}}(\mathcal{D}^{k}_{\text{loc}})$
7
8
9      $\theta_{\text{GAT}}(c)\leftarrow\sum_{k=1}^{K}\theta_{\text{GAT}}^{k}(N_{l})$triangleright server aggregation
10
return ${\theta}^{*}_{\text{GAT}}:=\theta_{\text{GAT}}(N_{c})$
Algorithm 1 Training of the GAT prior
3.2.2 MR Image Synthesis with the GAT Prior
To synthesize multi-site MR images via the GAT prior, the autoregressive transformer and VAE decoder modules are utilized. The transformer module generates a random set of multi-scale discrete token maps, sequentially across growing spatial scales based on the autoregressive mapping in Eq. 20. To introduce stochasticity during inference, however, nucleus sampling [62] is employed instead of the maximum-probability sampling procedure in Eqs. 18-19. Accordingly, the logit values across $V$ codebook vectors are ordered separately for each token in $\hat{\mathbf{f}}_{s}$, the top-5 percentile of values are subjected to softmax to obtain a probability mass function, and the predicted token is randomly drawn from this function:
[…]
$\tilde{\mathbf{P}}_{s}(i,:)$ $=$ $\mathrm{Softmax}(\mathbf{logit}_{\mathrm{sort}}(i,1:\lceil 0.05\,V\rceil)),$ (22)
[…]
Once a set of discrete multi-scale token maps $\{\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{S}\}$ is generated, it can be projected through the VAE decoder as described in Eq. 10 to obtain a synthetic MR image. For brevity, here we refer to the overall mapping used to synthesize images from site $k$ as follows:
$\mathbf{x}^{k}_{\text{syn}}$ $=\mathrm{GAT}(\mathbf{sp}(k)),\mbox{ such that:}$: (24)
$\mathrm{GAT}(\cdot)$ $=\mathrm{Dec}\left(\left\{(\mathbf{f}_{1},\dots,\mathbf{f}_{S}):\,\right.\right.$
$\qquad\quad\left.\left.\mathbf{f}_{s}=\mathrm{Trans}(\cdot,\mathbf{f}_{<s})\mbox{ for }s\in[1\mbox{ }S];\right\}\right).$
3.2.3 Two-Tier Training Procedure for FedGAT
To support model-heterogeneous settings, FedGAT leverages a two-tier training procedure where building of the global GAT prior (Alg. 1) is decoupled from subsequent building of site-specific reconstruction models (Alg. 2). The procedural details and loss functions used in each tier are discussed below.
Training of the GAT prior: FedGAT decentrally trains a global, site-prompted GAT prior that captures the distribution of multi-site MR images, which is later used to generate synthetic MR images from desired sites. The proposed GAT prior embodies VAE and autoregressive transformer modules as described in Section 3.2.1. VAE aims to encode input MR images onto a compact latent space and then to decode them back faithfully. Accordingly, a composite loss function is used for the VAE modules [59]:
$\mathcal{L}^{k}_{\text{vae}}(\mathcal{D}_{\text{loc}}^{k})=\mathbb{E}_{\mathbf{x}^{k}\sim\mathcal{D}_{\text{loc}}^{k}}\left[\|\mathbf{x}^{k}-\hat{\mathbf{x}}^{k}\|_{2}^{2}+\|\mathbf{z}^{k}-\hat{\mathbf{z}}^{k}\|_{2}^{2}\right.$
$\left.+\lambda_{p}\mathrm{LPIPS}(\mathbf{x}^{k},\hat{\mathbf{x}}^{k})+\lambda_{a}\mathcal{L}_{a}(\hat{\mathbf{x}}^{k})\right],$ (25)
where $k$ is the site index, and $\mathcal{D}_{\text{loc}}^{k}$ is the local training dataset. In Eq. 3.2.3, the first term enforces consistency between the input MR image and its prediction, the second term enforces consistency between the continuous representation in the latent space and its prediction, the third term is a perceptual loss [63], and the fourth term is an adversarial loss [64]. Meanwhile, the autoregressive transformer aims to autoregressively predict the discrete token maps across multiple scales faithfully. Since the discrete token maps are assigned categorically to individual vectors from the VAE codebook, the transformer module uses the following cross-entropy loss:
$\mathcal{L}^{k}_{\text{trans}}(\mathcal{D}_{\text{loc}}^{k})=-\mathbb{E}_{\mathbf{x}^{k}\sim\mathcal{D}_{\text{loc}}^{k}}\left[\sum_{s=1}^{S}\sum_{i=1}^{h_{s}w_{s}}\sum_{v=1}^{V}\mathbb{I}(\mathbf{f}_{s}^{k}(i)=v)\right.$
$\cdot\log\left(\mathbf{P}_{s}(i,v)\right)\Bigg{]},$ (26)
where the ground-truth discrete token maps are extracted by the VAE as $\{\mathbf{f}_{1}^{k},\mathbf{f}_{2}^{k},\dots,\mathbf{f}_{S}^{k}\}=\mathrm{Enc}(\mathbf{x}^{k})$, $i$ denotes token index within $\mathbf{f}^{k}_{s}$, $\mathbb{I}$ is a binary indicator that signals whether $\mathbf{f}^{k}_{s}(i)$ is categorized as the $v$th vector in the VAE codebook. Finally, the overall loss for the GAT prior can be expressed as:
$\mathcal{L}^{k}_{\text{GAT}}(\mathcal{D}_{\text{loc}}^{k})=\mathcal{L}^{k}_{\text{trans}}(\mathcal{D}_{\text{loc}}^{k})+\lambda_{\text{vae}}\mathcal{L}^{k}_{\text{vae}}(\mathcal{D}_{\text{loc}}^{k}),$ (27)
where $\lambda_{\text{vae}}$ denotes the relative weighting of the VAE loss.
As the architecture of the GAT prior is common across sites, decentralized training of the global prior across $K$ sites can be performed using the local loss function in Eq. 27 and conventional federated averaging of model weights based on a server-client topology as outlined in Alg. 1. The FL server coordinates decentralized training over $N_{c}$ communication rounds. At the start of the $c$th round, the global prior $\theta_{\text{GAT}}$ is sent to individual sites to initialize their local copies $\theta^{k}_{\text{GAT}}$:
$\theta^{k}_{\text{GAT}}(0)\leftarrow\theta_{\text{GAT}}(c-1),\,\,\forall k\in\{1,\ldots,K\}.$ (28)
Each site then updates its local copy by training it on its local dataset over $N_{l}$ epochs:
$\theta_{\text{GAT}}^{k}(e)\leftarrow\theta_{\text{GAT}}^{k}(e-1)-\eta\nabla_{\theta_{\text{GAT}}^{k}}\mathcal{L}^{k}_{\text{GAT}},\,\,\mbox{for }e\in[1\mbox{ }N_{l}],$ (29)
where $\eta$ denotes the learning rate. At the end of the round, local copies are forwarded to the server that aggregates them into the global prior via simple averaging assuming balanced numbers of samples across sites [36]:
$\theta_{\text{GAT}}(c)\leftarrow\sum_{k=1}^{K}\theta^{k}_{\text{GAT}}(N_{l}).$ (30)
At the completion of the decentralized training procedure, the trained global prior is taken as ${\theta}^{*}_{\text{GAT}}:=\theta_{\text{GAT}}(N_{c})$.
Input: $\mathcal{D}^{k}_{\text{\text{loc}}}$: local dataset from site $k$.
GAT: global GAT prior with ${\theta}^{*}_{\text{GAT}}$.
$N_{p}$: number of pre-training epochs.
$N_{f}$: number of fine-tuning epochs.
Output: $H^{k}_{{\phi}^{k}}$: reconstruction model at site $k$.
1
2Initialize model with $\phi^{k}(0)$ at site $k$.
3 $\triangleright$ pre-train on local dataset
4 for $e=1$ to $N_{p}$ do
5       $\phi^{k}(e)\leftarrow\phi^{k}(e-1)-\eta\nabla_{\phi^{k}}\mathcal{L}^{k}_{\text{rec}}(\mathcal{D}^{k}_{\text{\text{loc}}})$
6$\triangleright$ generate synthetic datasets
7 for $j=1$ to $K$ do
8       $\mathcal{D}^{j}_{\text{syn}}=\{\mathbf{x}^{j}_{\text{syn}},\mathcal{A}^{k}\mathbf{x}^{j}_{\text{syn}}\}$ where $\mathbf{x}^{j}_{\text{syn}}=\text{GAT}(\mathbf{sp}(j))$
9
10$\triangleright$ fine-tune on local & synthetic datasets
11 for $e=(1+N_{p})$ to $(N_{f}+N_{p})$ do
[…]
return ${\phi}^{*k}:={\phi}^{k}(N_{f}+N_{p})$
Algorithm 2 Training of reconstruction models
Training of reconstruction models: Following the first tier, each individual site trains a site-specific reconstruction model $H^{k}_{\phi^{k}}$ of preferred architecture, with the aid of the global GAT prior (Alg. 2). To attain a refined balance between within-site and cross-site reconstruction performances, the reconstruction model is built by pre-training on the local dataset followed by fine-tuning on a hybrid dataset containing both local and synthetic datasets. For pre-training, we employ a conventional mean-squared error loss [14]:
$\mathcal{L}^{k}_{\text{rec}}(\mathcal{D}_{\text{loc}}^{k})=\mathbb{E}_{(\mathbf{x}^{k},\mathbf{y}^{k})\sim\mathcal{D}_{\text{loc}}^{k}}\left[\|\mathbf{x}^{k}-H^{k}_{\phi^{k}}(\mathcal{F}^{-1}\mathbf{y}^{k},\mathcal{A}^{k})\|_{2}^{2}\right],$ (31)
where $\mathbf{x}^{k}$ is a reference image and $\mathbf{y}^{k}$ is the respective undersampled acquisition drawn from $\mathcal{D}_{\text{loc}}^{k}$. Pre-trained models are obtained after $N_{p}$ epochs:
$\phi^{k}(e)\leftarrow\phi^{k}(e-1)-\eta\nabla_{\phi^{k}}\mathcal{L}^{k}_{\text{rec}}(\mathcal{D}^{k}_{\text{\text{loc}}}),\mbox{ for }e\in[1\mbox{ }N_{p}].$ (32)
Next, each site generates synthetic coil-combined MR images from remaining sites via the site-prompted GAT prior to construct the hybrid dataset for fine-tuning. Note that training of reconstruction models does not only require coil-combined MR images but also respective multi-coil undersampled k-space data. To generate synthetic multi-coil k-space data, here we randomly couple synthetic coil-combined MR images with actual imaging operators derived from samples in the local dataset $\mathcal{D}^{k}_{\text{loc}}$ [10, 7]. In particular, for each sample in $\mathcal{D}_{\text{loc}}^{k}$, $\mathcal{A}^{k}=\mathcal{M}^{k}\mathcal{F}\mathcal{C}^{k}$ is constructed by conjoining coil sensitivity estimates derived from the fully-sampled acquisition ($\mathcal{C}^{k}$) with a random k-space sampling pattern ($\mathcal{M}^{k}$) and Fourier transformation ($\mathcal{F}$). Thus, synthetic MR images and multi-coil k-space data at site $k$ are derived as:
$\mathcal{D}^{j}_{\text{syn}}=\{\mathbf{x}^{j}_{\text{syn}},\mathbf{y}^{j}_{\text{syn}}\}\quad\forall j\in\{1,..,K\}\setminus\{k\};$ (33)
[…]
$\qquad\qquad\,\,\,\,\mathbf{y}^{j}_{\text{syn}}=\mathcal{A}^{k}\mathbf{x}^{j}_{\text{syn}}.$ (35)
The local MRI dataset and synthetic MRI datasets from remaining sites are then mixed in equal proportions of samples per site in order to construct a hybrid training set:
$\mathcal{D}^{k}_{\text{\text{hyb}}}=[\mathcal{D}^{k}_{\text{\text{loc}}};\,\mathcal{D}_{\text{syn}}^{\{1,..,K\}\setminus\{k\}}].$ (36)
Afterwards, the pre-trained reconstruction model is fine-tuned via mean-squared error loss:
$\mathcal{L}^{k}_{\text{rec}}(\mathcal{D}^{k}_{\text{\text{hyb}}})=\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim\mathcal{D}^{k}_{\text{hyb}}}\left[\|\mathbf{x}-H^{k}_{\phi^{k}}(\mathcal{F}^{-1}\mathbf{y},\mathcal{A}^{k})\|_{2}^{2}\right],$ (37)
where $\mathbf{x}$ is a reference image and $\mathbf{y}$ is the respective undersampled acquisition drawn from $\mathcal{D}_{\text{hyb}}^{k}$. Accordingly, fine-tuned models are obtained after $N_{f}$ epochs:
$\phi^{k}(e)\leftarrow\phi^{k}(e-1)-\eta\nabla_{\phi^{k}}\mathcal{L}^{k}_{\text{rec}}(\mathcal{D}^{k}_{\text{\text{hyb}}}),\mbox{ for }e\in[1+N_{p}\mbox{ }N_{f}+N_{p}].$ (38)
The model at the end of fine-tuning is taken as the final site-specific reconstruction model $\phi^{*k}:=\phi^{k}(N_{f}+N_{p})$.
4 Methods
4.1 Implementation Details for FedGAT
The proposed GAT prior embodies a VAE encoder to map MR images onto multi-scale discrete token maps [56], a transformer module to autoregressively generate discrete token maps across spatial scales by starting from a random embedding at the lowest scale [65], and a VAE decoder to recover images from discrete token maps at the highest scale. Here, GAT was implemented with $S$=\,=\,=10 scales of spatial dimensions taken as $h_{s}$=\,=\,=$w_{s}$=\,=\,=$p$, where $p$ varied from lowest to highest scale as $\{1,2,3,4,5,6,8,10,13,16\}$. The VAE encoder and decoder each had 6 stages. Across stages, the encoder used a total of 12 residual convolutional blocks [58], and performed spatial downsampling of the input MR image onto the continuous latent representation via a set of stride-2 convolutions. The decoder used a total of 12 residual convolutional blocks, and performed spatial upsampling of the continuous latent representation via transpose convolutions. A channel dimensionality of $c$=\,=\,=32, and a codebook size of $V$=\,=\,=4096 was prescribed. The autoregressive transformer module employed $L$=\,=\,=16 sequential blocks with embedding dimensionality of $d$=\,=\,=1024 and 16 self-attention heads.
For the first tier of the training procedure, $N_{c}$=\,=\,=500 communication rounds and $N_{l}$=\,=\,=1 local epochs per round were used. To facilitate federated learning of the GAT prior, here the number of trained and thereby communicated model parameters were reduced by using frozen VAE modules with weights initialized from the ‘vae_ch160v4096z32’ architecture ( [^1] https://huggingface.co/FoundationVision/var/resolve/main/vae_ch160v4096z32.pthpre-trained on natural images with $\lambda_{p}=1$, $\lambda_{a}=0.4$), and setting $\lambda_{vae}$=\,=\,=0. Meanwhile, the weights of the transformer module were randomly initialized. For the second tier, $N_{p}$=\,=\,=100 pre-training epochs and $N_{f}$=\,=\,=100 fine-tuning epochs were used.
4.2 Competing Methods
FedGAT was comparatively demonstrated against non-FL benchmarks (Central, Single) along with distillation-based (FedMD, FedDF) and generative (FedGIMP, FedDDA) FL baselines.
Central: A non-FL benchmark for across-site reconstruction performance was considered that pools MRI datasets across sites to centrally train a global reconstruction model [25]. Since Central requires selection of a homogeneous architecture across sites, in model-heterogeneous settings, separate global reconstruction models were trained based on $K$ distinct architectures preferred by different sites.
Single: A non-FL benchmark for within-site reconstruction performance was considered that trains site-specific reconstruction models on local data [30]. Since reconstruction models are trained independently across sites, Single is natively compatible with model-heterogeneous settings.
FedDF: A distillation-based FL baseline was considered that supports model-heterogeneous settings via an auxiliary public dataset [66]. To adopt FedDF for MRI reconstruction, distillation between site-specific reconstruction models and a global model were performed based on predictions of reconstructed images, instead of class logits as in the original implementation for classification tasks [66]. At individual sites, site-specific reconstruction models were initialized via distillation from the global model (taken as MoDL-5), and then trained along with the copy of the global model using a proximal loss term [67]. Training was performed on a 840-sample mixture of the local data with independent auxiliary data downloaded from a public repository [68]. On the server, copies of the global model locally trained at individual sites were aggregated via distillation on a 840-sample set of auxiliary data.
FedMD: A distillation-based FL baseline was considered that supports model-heterogeneous settings via an auxiliary public dataset [69]. To adopt FedMD for MRI reconstruction, predictions of reconstructed images were used for knowledge transfer instead of class logits in the original implementation for classification tasks [69]. Accordingly, site-specific reconstruction models were locally trained on a 840-sample mixture of local and auxiliary data as in FedDF [68], images reconstructed by site-specific models on the auxiliary dataset were aggregated on the server [36], and the aggregated reconstructions on the auxiliary dataset were used to initialize site-specific models via distillation.
FedGIMP: A generative FL baseline that decentrally trains a global prior based on an unconditional adversarial model to synthesize multi-site MR images [31]. Note that FedGIMP was originally devised for model-homogeneous settings. To adopt FedGIMP for model-heterogeneous settings, it was implemented using the same two-tier training strategy as in FedGAT, where site-specific reconstruction models were trained using a hybrid dataset containing both local MRI data and synthetic data from remaining sites generated by FedGIMP.
FedDDA: A generative FL baseline that decentrally trains a global prior based on an unconditional diffusion model to synthesize images [50]. A site index was included in FedDDA for controllable generation of multi-site MR images [42]. FedDDA was imeplemented using the same two-tier training strategy as in FedGAT to support model-heterogeneous settings, where site-specific reconstruction models were trained using a hybrid dataset containing both local MRI data and synthetic data from remaining sites generated by FedDDA.
Table 1: Reconstruction performance for single-coil datasets. Each site employs an entirely distinct model type, with Site $k$ (dataset-type) denoting the site index, the local dataset, and the type of reconstruction model. Metrics are listed as mean±$\pm$±std across the test set, for within-site (upper panel) and across-site (lower panel) reconstructions at R=4x-8x. Boldface indicates the top-performing FL method in each task. Benchmark methods for within-site reconstruction (Single), and for cross-site reconstruction (Central) are underlined.
Within-site Site 1 (BraTS-MoDL) Site 2 (fastMRI-rGAN) Site 3 (MIDAS-D5C5)
reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
non-FL Single 4x 48.06$\pm$3.33 99.72$\pm$0.11 35.83$\pm$2.51 90.64$\pm$3.87 29.04$\pm$1.84 76.77$\pm$4.08
8x 42.46$\pm$3.15 99.03$\pm$0.45 33.18$\pm$2.26 86.58$\pm$4.64 27.79$\pm$2.05 74.20$\pm$3.97
Central 4x 46.12$\pm$2.55 99.46$\pm$0.20 36.03$\pm$2.51 91.23$\pm$3.63 28.89$\pm$1.89 75.11$\pm$4.07
8x 42.16$\pm$2.72 98.65$\pm$0.58 33.18$\pm$2.53 87.40$\pm$4.44 27.34$\pm$2.17 73.87$\pm$3.98
FL methods FedDF 4x 44.42$\pm$3.76 99.34$\pm$0.28 32.88$\pm$3.22 87.95$\pm$4.41 27.62$\pm$1.90 61.23$\pm$4.32
8x 40.06$\pm$3.23 98.23$\pm$0.77 30.78$\pm$2.73 83.45$\pm$5.51 27.48$\pm$1.93 59.62$\pm$4.93
FedMD 4x 44.68$\pm$3.72 99.40$\pm$0.24 34.24$\pm$3.06 88.48$\pm$4.13 28.95$\pm$1.92 62.22$\pm$5.25
8x 40.32$\pm$3.22 98.29$\pm$0.75 32.96$\pm$2.56 86.40$\pm$4.66 27.79$\pm$2.07 60.19$\pm$5.07
FedGIMP 4x 47.61$\pm$2.86 99.63$\pm$0.15 35.45$\pm$2.48 90.27$\pm$3.99 28.68$\pm$1.90 74.91$\pm$4.42
8x 42.14$\pm$2.71 98.62$\pm$0.54 32.74$\pm$2.40 85.86$\pm$4.75 26.82$\pm$2.17 73.96$\pm$4.08
FedDDA 4x 47.50$\pm$2.81 99.61$\pm$0.15 35.67$\pm$2.51 90.30$\pm$3.94 28.61$\pm$1.79 75.67$\pm$4.15
8x 41.76$\pm$2.90 98.47$\pm$0.63 32.79$\pm$2.45 86.02$\pm$4.84 27.10$\pm$2.18 73.08$\pm$4.07
FedGAT 4x 47.63$\pm$2.65 99.64$\pm$0.14 35.89$\pm$2.50 90.69$\pm$3.86 31.88$\pm$2.03 77.48$\pm$4.94
8x 41.60$\pm$2.85 98.78$\pm$0.54 33.23$\pm$2.36 86.45$\pm$0.67 29.62$\pm$2.11 74.88$\pm$4.29
Across-site Site 1 (BraTS-MoDL) Site 2 (fastMRI-rGAN) Site 3 (MIDAS-D5C5)
reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
non-FL Single 4x 34.45$\pm$3.57 91.44$\pm$4.05 33.02$\pm$2.30 66.49$\pm$7.26 29.67$\pm$2.28 75.67$\pm$6.93
8x 30.48$\pm$3.20 85.48$\pm$5.37 30.08$\pm$2.20 59.14$\pm$6.78 28.38$\pm$2.26 73.64$\pm$6.55
Central 4x 36.48$\pm$2.89 94.78$\pm$1.91 33.53$\pm$2.31 82.93$\pm$4.07 30.68$\pm$2.47 77.41$\pm$6.99
8x 32.79$\pm$2.81 92.16$\pm$2.62 30.78$\pm$2.25 78.75$\pm$4.67 29.93$\pm$2.20 75.86$\pm$6.95
FL methods FedDF 4x 34.40$\pm$3.12 93.12$\pm$3.00 33.14$\pm$2.61 71.59$\pm$6.74 29.23$\pm$2.38 72.76$\pm$7.12
8x 30.91$\pm$2.87 90.33$\pm$3.36 30.89$\pm$2.22 64.17$\pm$7.14 28.37$\pm$2.63 72.45$\pm$8.43
FedMD 4x 34.61$\pm$3.13 93.42$\pm$2.75 33.11$\pm$2.26 68.04$\pm$6.84 30.03$\pm$2.41 72.31$\pm$7.93
8x 31.12$\pm$2.83 90.63$\pm$3.13 30.47$\pm$2.20 61.79$\pm$7.06 29.05$\pm$2.57 71.80$\pm$7.84
FedGIMP 4x 36.42$\pm$3.17 94.32$\pm$2.49 33.05$\pm$2.29 67.50$\pm$7.01 30.30$\pm$2.33 76.19$\pm$6.56
8x 31.88$\pm$2.87 90.16$\pm$3.78 29.78$\pm$2.16 57.80$\pm$6.39 28.31$\pm$2.16 71.75$\pm$6.10
FedDDA 4x 36.19$\pm$3.13 94.41$\pm$2.24 33.11$\pm$2.21 67.78$\pm$7.00 28.59$\pm$2.07 78.13$\pm$5.12
8x 31.79$\pm$2.86 90.75$\pm$3.07 30.15$\pm$2.16 59.18$\pm$6.93 27.73$\pm$2.13 77.92$\pm$5.11
FedGAT 4x 36.71$\pm$3.01 95.02$\pm$1.96 34.50$\pm$2.36 75.69$\pm$5.76 30.77$\pm$2.35 78.62$\pm$6.41
8x 32.02$\pm$2.95 91.53$\pm$3.07 31.47$\pm$2.49 71.84$\pm$6.47 29.72$\pm$2.38 78.08$\pm$6.86
Table 2: Within-site and cross-site reconstruction performance for multi-coil datasets (f.knee denotes fastMRI-knee, f.brain denotes fastMRI-brain, UMRAM denotes the in-house dataset), where each site employs an entirely different model type.
Within-site Site 1 (f.knee-MoDL) Site 2 (f.brain-rGAN) Site 3 (UMRAM-D5C5)
reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
non-FL Single 4x 33.78$\pm$2.27 93.37$\pm$1.52 36.97$\pm$3.07 97.32$\pm$1.17 31.81$\pm$1.73 90.10$\pm$2.57
8x 29.39$\pm$2.06 83.90$\pm$3.08 33.08$\pm$3.39 94.54$\pm$2.26 28.77$\pm$1.71 85.29$\pm$3.64
Central 4x 33.69$\pm$2.32 93.20$\pm$1.63 36.93$\pm$3.22 97.32$\pm$1.19 31.02$\pm$1.60 88.49$\pm$2.56
8x 29.39$\pm$2.19 83.62$\pm$3.09 33.79$\pm$3.05 94.59$\pm$2.13 27.39$\pm$1.68 82.72$\pm$3.04
FL methods FedDF 4x 32.79$\pm$2.19 90.40$\pm$2.45 33.95$\pm$2.88 96.30$\pm$1.39 30.34$\pm$1.55 88.99$\pm$2.53
8x 27.70$\pm$2.33 78.94$\pm$4.57 28.80$\pm$2.82 90.60$\pm$3.14 27.05$\pm$1.72 83.94$\pm$3.40
FedMD 4x 32.48$\pm$2.24 90.57$\pm$2.21 34.07$\pm$2.83 96.36$\pm$1.36 30.10$\pm$1.61 88.73$\pm$2.52
8x 26.99$\pm$2.47 78.49$\pm$4.51 29.08$\pm$3.15 90.16$\pm$3.23 27.13$\pm$1.64 84.18$\pm$3.31
FedGIMP 4x 33.78$\pm$2.27 93.41$\pm$1.48 31.72$\pm$3.85 94.25$\pm$4.32 30.11$\pm$2.08 88.56$\pm$2.43
8x 28.97$\pm$2.12 83.68$\pm$3.07 28.78$\pm$3.10 90.71$\pm$5.21 26.26$\pm$2.68 82.69$\pm$3.52
FedDDA 4x 32.53$\pm$2.33 91.75$\pm$1.82 32.69$\pm$3.46 94.71$\pm$3.44 24.47$\pm$2.57 81.20$\pm$3.47
8x 28.09$\pm$2.22 81.90$\pm$2.93 29.63$\pm$3.28 90.85$\pm$5.46 18.38$\pm$1.97 68.73$\pm$5.73
FedGAT 4x 34.10$\pm$2.24 97.32$\pm$1.67 35.86$\pm$2.78 96.85$\pm$1.19 31.94$\pm$1.70 90.21$\pm$2.51
8x 29.49$\pm$2.21 83.73$\pm$3.07 31.48$\pm$2.34 92.89$\pm$2.38 28.28$\pm$1.89 85.20$\pm$3.58
Across-site Site 1 (f.knee-MoDL) Site 2 (f.brain-rGAN) Site 3 (UMRAM-D5C5)
reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
non-FL Single 4x 38.40$\pm$3.26 97.32$\pm$1.67 31.17$\pm$3.22 93.03$\pm$2.57 26.18$\pm$2.62 79.42$\pm$5.73
8x 32.39$\pm$3.04 92.63$\pm$2.94 26.82$\pm$3.15 85.13$\pm$5.06 22.64$\pm$2.30 70.31$\pm$6.85
Central 4x 39.70$\pm$3.36 97.62$\pm$1.70 37.41$\pm$2.13 96.27$\pm$1.29 27.52$\pm$2.49 81.63$\pm$5.35
8x 34.33$\pm$3.12 93.86$\pm$3.04 33.98$\pm$2.11 91.89$\pm$2.82 25.09$\pm$2.47 74.38$\pm$6.46
FL methods FedDF 4x 38.03$\pm$2.99 96.54$\pm$1.41 32.30$\pm$2.93 92.69$\pm$1.76 26.68$\pm$2.55 81.09$\pm$5.31
8x 32.05$\pm$2.80 92.25$\pm$3.12 27.68$\pm$2.80 85.12$\pm$4.50 24.04$\pm$2.48 70.60$\pm$6.94
FedMD 4x 37.77$\pm$3.07 96.53$\pm$1.59 32.62$\pm$2.94 92.69$\pm$1.69 26.39$\pm$2.50 80.99$\pm$5.26
8x 31.35$\pm$2.91 91.22$\pm$3.24 28.03$\pm$2.50 84.44$\pm$4.68 23.99$\pm$2.43 70.13$\pm$6.61
FedGIMP 4x 37.87$\pm$3.13 97.23$\pm$1.50 26.57$\pm$3.13 85.71$\pm$5.69 26.00$\pm$2.49 80.03$\pm$5.68
8x 32.01$\pm$2.98 92.52$\pm$2.91 22.85$\pm$2.93 75.96$\pm$8.28 22.20$\pm$2.29 67.89$\pm$8.82
FedDDA 4x 35.54$\pm$3.34 94.49$\pm$2.44 27.04$\pm$3.51 84.62$\pm$4.68 22.34$\pm$2.33 72.91$\pm$6.55
8x 30.15$\pm$3.57 88.74$\pm$3.98 23.98$\pm$3.40 75.66$\pm$5.86 19.21$\pm$2.32 60.73$\pm$9.18
FedGAT 4x 39.25$\pm$3.09 97.50$\pm$1.52 33.95$\pm$2.34 93.45$\pm$1.95 27.17$\pm$2.58 81.24$\pm$5.63
8x 33.52$\pm$2.98 92.90$\pm$3.11 29.20$\pm$2.28 86.58$\pm$4.23 24.45$\pm$2.55 77.63$\pm$8.77
4.3 Modeling Procedures
We considered several three-site FL setups to examine performance under model-heterogeneous settings. The reconstruction models trained in these setups included MoDL [14], rGAN [20], and D5C5 [13]. All sites were taken to use common imaging operators with matched acceleration rates and k-space sampling densities. Each site was allowed to train a distinct reconstruction model. In a first experiment, the three sites trained entirely distinct architectures (MoDL-3, rGAN, D5C5). In a second experiment, the sites trained the same model type albeit using different numbers of cascades (MoDL-3, MoDL-5, MoDL-7). All models were implemented using the PyTorch framework and executed on RTX 4090 GPUs. Models were trained via the AdamW optimizer with $\beta_{1}$=\,=\,=0.9, $\beta_{2}$=\,=\,=0.95, a weight decay of 0.05, and a learning rate of $\eta$=\,=\,=$10^{-4}$.
As model-heterogeneous settings preclude aggregation via simple averaging of model weights, FL baselines used either model distillation or synthetic image generation as a means for cross-site knowledge transfer. Note that distillation-based aggregation involves bidirectional distillation between global and site-specific models, with each direction incurring significantly higher computational costs compared to model averaging, as the student model must be trained based on a sizable set of input-output images produced by the teacher model. Here, we observed that distillation-based baselines (FedDF, FedMD) attained a near-optimal trade-off between performance and efficiency with a total of 200 training epochs and distillation every 100 epochs. In generative baselines (FedGIMP, FedDDA), the two-tier strategy in FedGAT was followed under matching training procedures with $N_{c}$=\,=\,=500, $N_{l}$=\,=\,=1, $N_{p}$=\,=\,=100 and $N_{f}$=\,=\,=100.
Reconstruction performance was evaluated via peak signal-to-noise ratio (PSNR) and structural similarity (SSIM) metrics. Fidelity of synthetic MR images was evaluated via Freschet’s inception distance (FID) metric. Performance differences among competing methods were examined via non-parametric Wilcoxon signed-rank tests (p$<$0.05).
Figure 3: Representative reconstructions at R=8x from zero-filled Fourier method (Zero-filled), single-site models (Single), FL baselines (FedDF, FedMD, FedGIMP, FedDDA), and FedGAT, along with reference images. (a) Site-specific BraTS model tested on fastMRI, (b) Site-specific fastMRI model tested on MIDAS, (c) Site-specific MIDAS model on BraTS. Zoom-in displays are included showing error maps (left) and images (right) within a focal window to emphasize method differences.
Table 3: Within-site and cross-site reconstruction performance for single-coil datasets, where each site employs a variant of MoDL with different number of cascades.
Within-site Site 1 (BraTS-MoDL-5) Site 2 (fastMRI-MoDL-3) Site 3 (MIDAS-MoDL-7)
Reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
Non-FL Single 4x 51.80$\pm$2.83 99.86$\pm$0.06 40.79$\pm$2.94 97.09$\pm$1.26 36.31$\pm$2.89 95.91$\pm$1.71
8x 44.07$\pm$3.26 99.38$\pm$0.33 36.30$\pm$2.83 95.14$\pm$1.81 30.93$\pm$2.81 92.45$\pm$3.01
Central 4x 51.30$\pm$2.38 99.83$\pm$0.08 40.00$\pm$3.12 96.76$\pm$1.36 36.06$\pm$2.76 95.73$\pm$1.73
8x 44.19$\pm$2.79 99.26$\pm$0.36 35.02$\pm$3.00 94.14$\pm$2.05 31.08$\pm$2.80 92.47$\pm$2.97
FL methods FedDF 4x 47.08$\pm$2.88 99.63$\pm$0.13 39.31$\pm$3.12 96.35$\pm$1.55 33.31$\pm$2.71 94.19$\pm$2.13
8x 41.53$\pm$3.13 98.89$\pm$0.46 34.18$\pm$3.13 93.49$\pm$2.48 29.54$\pm$2.55 90.28$\pm$3.28
FedMD 4x 47.02$\pm$2.84 99.63$\pm$0.13 39.23$\pm$3.27 96.37$\pm$1.58 33.37$\pm$2.69 94.20$\pm$2.12
8x 41.43$\pm$3.10 98.86$\pm$0.47 33.97$\pm$3.16 93.52$\pm$2.46 30.03$\pm$2.53 90.52$\pm$3.25
FedGIMP 4x 46.33$\pm$3.07 99.56$\pm$0.14 39.46$\pm$3.60 96.38$\pm$1.78 34.71$\pm$2.88 94.66$\pm$2.21
8x 41.60$\pm$2.89 98.72$\pm$0.42 34.90$\pm$3.32 93.97$\pm$2.57 30.28$\pm$2.78 90.56$\pm$3.63
FedDDA 4x 47.38$\pm$2.81 99.66$\pm$0.12 39.22$\pm$3.35 96.53$\pm$1.51 29.16$\pm$1.39 60.74$\pm$2.14
8x 42.92$\pm$2.99 99.08$\pm$0.40 34.14$\pm$3.36 93.86$\pm$2.43 27.80$\pm$2.03 63.50$\pm$2.74
FedGAT 4x 50.42$\pm$2.44 99.81$\pm$0.08 40.01$\pm$3.20 96.78$\pm$1.35 35.33$\pm$2.90 95.57$\pm$1.72
8x 43.43$\pm$2.85 99.25$\pm$0.35 35.50$\pm$2.91 94.54$\pm$2.10 30.04$\pm$2.87 91.92$\pm$3.10
Across-site Site 1 (BraTS-MoDL-5) Site 2 (fastMRI-MoDL-3) Site 3 (MIDAS-MoDL-7)
Reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
Non-FL Single 4x 34.21$\pm$4.48 90.92$\pm$4.36 39.25$\pm$6.94 95.83$\pm$4.14 45.38$\pm$5.04 98.28$\pm$1.77
8x 30.79$\pm$4.08 87.83$\pm$5.52 34.46$\pm$6.48 92.76$\pm$5.57 38.99$\pm$4.18 97.09$\pm$2.44
Central 4x 38.98$\pm$4.55 96.36$\pm$1.80 40.83$\pm$7.05 96.78$\pm$3.19 47.23$\pm$5.35 98.62$\pm$1.50
8x 33.82$\pm$4.45 93.52$\pm$3.15 35.31$\pm$6.20 93.43$\pm$5.26 41.77$\pm$4.41 97.75$\pm$2.03
FL methods FedDF 4x 36.36$\pm$2.96 95.21$\pm$1.92 39.33$\pm$2.66 96.23$\pm$1.74 43.44$\pm$3.03 98.09$\pm$1.04
8x 31.94$\pm$2.87 92.09$\pm$2.86 34.91$\pm$2.80 93.62$\pm$2.48 37.86$\pm$3.15 96.31$\pm$1.75
FedMD 4x 36.28$\pm$2.92 95.24$\pm$1.87 39.55$\pm$2.75 96.28$\pm$1.79 43.43$\pm$3.02 98.08$\pm$1.04
8x 31.91$\pm$2.85 92.16$\pm$2.81 34.87$\pm$2.84 93.63$\pm$2.53 37.58$\pm$3.25 96.37$\pm$1.72
FedGIMP 4x 36.97$\pm$4.56 94.87$\pm$3.09 38.71$\pm$6.51 95.25$\pm$4.62 44.10$\pm$4.08 98.33$\pm$1.74
8x 32.89$\pm$4.50 92.33$\pm$3.89 34.65$\pm$6.20 92.31$\pm$5.76 39.02$\pm$3.74 96.84$\pm$2.79
FedDDA 4x 37.52$\pm$4.02 95.48$\pm$2.51 39.81$\pm$6.80 95.91$\pm$4.06 32.27$\pm$1.40 59.96$\pm$6.01
8x 32.98$\pm$4.21 92.71$\pm$3.53 35.06$\pm$6.15 93.16$\pm$5.36 31.45$\pm$1.78 59.85$\pm$7.96
FedGAT 4x 38.07$\pm$4.48 96.09$\pm$1.93 40.32$\pm$7.04 96.33$\pm$3.66 46.68$\pm$5.21 98.68$\pm$1.65
8x 33.35$\pm$4.23 93.16$\pm$3.37 35.13$\pm$6.94 93.64$\pm$5.43 40.65$\pm$4.37 97.42$\pm$2.31
Table 4: Within-site and cross-site reconstruction performance for multi-coil datasets, where each site employs a variant of MoDL with different number of cascades.
Within-site Site 1 (f.knee-MoDL-3) Site 2 (f.brain-MoDL-5) Site 3 (UMRAM-MoDL-7)
Reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
Non-FL Single 4x 33.78$\pm$2.27 93.37$\pm$1.52 40.33$\pm$3.45 98.03$\pm$0.84 49.96$\pm$2.06 99.61$\pm$0.17
8x 29.39$\pm$2.06 83.90$\pm$3.08 34.21$\pm$3.36 94.33$\pm$2.01 42.11$\pm$1.90 98.25$\pm$0.73
Central 4x 33.69$\pm$2.32 93.20$\pm$1.63 40.95$\pm$3.84 98.05$\pm$1.19 46.36$\pm$2.22 99.32$\pm$0.30
8x 29.39$\pm$2.19 83.62$\pm$3.09 34.27$\pm$3.50 94.11$\pm$2.28 39.73$\pm$1.86 97.87$\pm$0.72
FL methods FedDF 4x 32.33$\pm$2.36 90.47$\pm$2.44 38.87$\pm$3.62 95.66$\pm$1.49 44.56$\pm$2.72 98.52$\pm$0.63
8x 28.36$\pm$2.31 80.66$\pm$4.14 32.97$\pm$3.73 91.37$\pm$3.37 37.00$\pm$2.30 96.73$\pm$1.40
FedMD 4x 32.93$\pm$2.35 91.96$\pm$1.77 39.94$\pm$3.94 97.49$\pm$1.09 44.63$\pm$3.05 99.25$\pm$0.28
8x 28.15$\pm$2.42 81.60$\pm$3.52 33.85$\pm$3.79 93.68$\pm$2.55 37.55$\pm$2.55 97.09$\pm$1.20
FedGIMP 4x 33.78$\pm$2.27 93.41$\pm$1.48 40.54$\pm$3.38 97.88$\pm$0.90 44.99$\pm$2.09 99.02$\pm$0.32
8x 28.97$\pm$2.12 83.68$\pm$3.07 34.19$\pm$3.37 93.80$\pm$3.22 39.53$\pm$1.70 97.68$\pm$0.77
FedDDA 4x 32.53$\pm$2.33 91.75$\pm$1.82 39.60$\pm$3.62 97.52$\pm$1.38 42.57$\pm$2.92 97.32$\pm$0.78
8x 28.09$\pm$2.22 81.90$\pm$2.93 33.25$\pm$3.52 93.61$\pm$2.76 37.51$\pm$2.17 93.63$\pm$1.50
FedGAT 4x 34.10$\pm$2.24 93.44$\pm$1.46 40.87$\pm$3.58 98.03$\pm$1.27 45.13$\pm$2.14 99.22$\pm$0.29
8x 29.49$\pm$2.21 83.73$\pm$3.07 34.20$\pm$3.48 93.85$\pm$2.52 40.21$\pm$1.71 97.84$\pm$0.66
Across-site Site 1 (f.knee-MoDL-3) Site 2 (f.brain-MoDL-5) Site 3 (UMRAM-MoDL-7)
Reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
Non-FL Single 4x 38.40$\pm$3.26 97.32$\pm$1.67 39.42$\pm$4.49 96.91$\pm$2.13 37.60$\pm$4.58 96.03$\pm$2.34
8x 32.39$\pm$3.04 92.63$\pm$2.94 32.99$\pm$3.57 90.55$\pm$6.10 31.48$\pm$3.96 88.07$\pm$5.35
Central 4x 39.70$\pm$3.36 97.62$\pm$1.70 40.64$\pm$4.84 97.36$\pm$1.99 40.39$\pm$4.11 97.53$\pm$1.31
8x 34.33$\pm$3.12 93.86$\pm$3.04 33.61$\pm$3.89 91.38$\pm$5.74 34.47$\pm$3.50 91.89$\pm$4.04
FL methods FedDF 4x 38.15$\pm$3.41 96.40$\pm$2.11 38.28$\pm$4.64 95.18$\pm$3.56 37.37$\pm$4.12 94.73$\pm$2.39
8x 32.94$\pm$3.12 92.43$\pm$3.47 32.69$\pm$4.11 89.39$\pm$7.34 30.15$\pm$3.64 86.84$\pm$5.61
FedMD 4x 38.75$\pm$3.44 97.01$\pm$1.71 39.32$\pm$4.78 96.80$\pm$2.28 38.39$\pm$4.49 95.92$\pm$2.28
8x 32.52$\pm$2.90 92.24$\pm$3.02 33.19$\pm$3.76 90.90$\pm$6.08 31.90$\pm$3.81 87.51$\pm$5.55
FedGIMP 4x 37.87$\pm$3.13 97.23$\pm$1.50 38.93$\pm$4.32 96.75$\pm$2.16 37.40$\pm$4.38 96.35$\pm$1.76
8x 32.01$\pm$2.98 92.52$\pm$2.91 32.96$\pm$3.63 90.71$\pm$5.94 31.70$\pm$4.13 89.29$\pm$5.18
FedDDA 4x 35.54$\pm$3.34 94.49$\pm$2.44 37.78$\pm$3.78 95.65$\pm$1.52 36.61$\pm$4.16 95.15$\pm$2.36
8x 30.15$\pm$3.57 88.74$\pm$3.98 32.10$\pm$3.43 89.21$\pm$5.03 30.80$\pm$3.74 86.32$\pm$5.53
FedGAT 4x 39.25$\pm$3.09 97.50$\pm$1.52 39.65$\pm$4.53 97.16$\pm$2.07 39.16$\pm$4.33 97.10$\pm$1.60
8x 33.52$\pm$2.98 92.90$\pm$3.11 33.58$\pm$3.93 90.93$\pm$5.81 32.98$\pm$3.89 90.29$\pm$4.57
4.4 Datasets
In single-coil experiments, multi-contrast coil-combined magnitude MR images from public BraTS [70], fastMRI [71] and MIDAS [72] datasets were analyzed. Each dataset was taken to correspond to a separate site. At each site, (training, validation, test) splits of (40, 5, 10) non-overlapping subjects were used, resulting in (840, 105, 210) cross-sectional images. Subjects were selected at random from the datasets.
In multi-coil experiments, multi-contrast k-space data from public fastMRI-brain, fastMRI-knee [71] and an in-house dataset were analyzed [31]. Experimental procedures for collecting the in-house dataset were approved by the local ethics committee at Bilkent University, and written informed consent was obtained from all participants. Each dataset was taken to correspond to a separate site. At each site, (training, validation, test) splits of (36, 6, 8) non-overlapping subjects were used, resulting in (840, 105, 210) cross-sectional images. Subjects were selected at random from the datasets. Multi-coil data were compressed to 5 virtual coils [73].
Both actual and synthetic k-space data were retrospectively undersampled using a variable-density (VD) pattern with acceleration factors R = 4x and 8x [1]. Models were trained on a mixture of multi-contrast MRI datasets without special procedures to handle separate contrasts. When necessary, actual and synthetic MR images were zero-padded to a consistent size to permit training of reconstruction models, albeit performance assessments were conducted in native image sizes by discarding the padded regions following reconstruction.
5 Results
5.1 Comparison Studies
To demonstrate the utility of FedGAT for enabling flexible collaborations in learning-based MRI reconstruction, we considered model-heterogeneous FL setups where each site used a different network architecture to build its reconstruction model. FedGAT was compared against several state-of-the-art baselines including distillation methods (FedDF, FedMD), and generative methods (FedGIMP, FedDDA). On one hand, site-specific performance was evaluated by conducting within-site reconstructions, where the eventual reconstruction model available at a given site was deployed on the same site (i.e., model for Site $k$ tested on data from Site $k$). ‘Single’ models trained on single-site data locally at each site were considered as a performance benchmark for these within-site reconstructions. On the other hand, generalization performance was evaluated by conducting across-site reconstructions, where the eventual reconstruction model available at a given site was deployed on remaining sites (i.e., model for Site $k$ tested on data from Site $l$ where $k\neq l$). ‘Central’ models trained on the server using multi-site data aggregated across all sites were taken as a privacy-violating performance benchmark for across-site reconstructions.
In a first FL setup, we examined performance when individual sites employed entirely distinct types of model architectures (i.e., MoDL-3, rGAN, D5C5). Performance metrics are listed in Table 1 for single-coil and in Table 2 for multi-coil MRI datasets. For within-site reconstructions, FedGAT achieves significantly higher performance than all competing FL methods (p$<$0.05), except for FedGIMP that yields higher PSNR for R=8x on single-coil data at Site 1. On average, FedGAT offers (PSNR, SSIM) improvements of (2.06dB, 4.40%) over distillation baselines, and (2.07dB, 2.48)SSIM over generative baselines. Among FL methods, FedGAT attains the most competitive performance against the ‘Single’ benchmark, and outperforms the privacy-violating ‘Central’ by 0.42dB PSNR, 0.71% SSIM (p$<$0.05). For across-site reconstructions, FedGAT achieves significantly higher performance than all competing FL methods (p$<$0.05). On average, FedGAT offers (PSNR, SSIM) improvements of (1.22dB, 3.48%) over distillation baselines, and (2.60dB, 5.74)SSIM over generative baselines. In this setting, FedGAT attains the most competitive performance against the ‘Central’ benchmark while outperforming ‘Single’ by 1.59dB PSNR, 4.20% SSIM (p$<$0.05), which showcasing the generalization benefits of training models on multi-site datasets. Representative images reconstructed via FL methods in the examined model-heterogeneous setup are depicted in Fig. 3. Corroborating the quantitative assessments, FedGAT attains higher image quality compared to baselines, with sharper depiction of structural details, lower artifact and noise in reconstructed images. Meanwhile, competing FL methods suffer from a degree of spatial blur, residual noise and artifacts that lead to suboptimal depiction of detailed tissue structures.
In a second FL setup, we examined performance when individual sites employed distinct variants of a common model type (i.e., MoDL with 3, 5, versus 7 cascades), resulting in varying levels of model complexity across sites. Naturally, the benefits of model-agnostic approaches can be lower in this setup, when compared against the first setup that prescribed entirely distinct model types. Still, we reasoned that FedGAT should continue to yield significant benefits over competing FL baselines. Performance metrics are listed in Table 3 for single-coil and in Table 4 for multi-coil MRI datasets. For within-site reconstructions, FedGAT achieves significantly higher performance than all competing FL methods, and attains the most competitive performances against the ‘Single’ and privacy-violating ‘Central’ (p$<$0.05), except for FedMD that yields higher PSNR on multi-coil data at Site 3. On average, FedGAT offers (PSNR, SSIM) improvements of (1.52dB, 1.15%) over distillation baselines, and (1.42dB, 3.32)SSIM over generative baselines. For across-site reconstructions, FedGAT achieves significantly higher performance than all competing FL methods (p$<$0.05). On average, FedGAT offers (PSNR, SSIM) improvements of (1.40dB, 0.99%) over distillation baselines, and (2.31dB, 4.29)SSIM over generative baselines. In this setting, FedGAT attains the most competitive performance against ‘Central’ while outperforming ‘Single’ by 1.42dB PSNR, 1.42% SSIM (p$<$0.05), indicating improved generalization. Representative images reconstructed in the second model-heterogeneous FL setup are depicted in Fig. 4. Corroborating the quantitative assessments, FedGAT attains higher image quality compared to baselines, with sharper depiction of structural details, lower artifacts and noise. Meanwhile, competing FL methods suffer from a degree of spatial blur, noise or artifacts that elicit degradations in depiction of detailed tissue structures.
Figure 4: Representative reconstructions at R=8x from zero-filled Fourier method (Zero-filled), single-site models (Single), FL baselines (FedDF, FedMD, FedGIMP, FedDDA), and FedGAT, along with reference images. (a) Site-specific fastMRI-knee model tested on UMRAM, (b) Site-specific fastMRI-brain model tested on fastMRI-knee, (c) Site-specific UMRAM model tested on fastMRI-brain. Zoom-in windows of error maps and images are included to emphasize method differences.
5.2 Ablation Studies
The primary aim of the current study is to enable collaborative training of generalizable reconstruction models with heterogeneous architectures across imaging sites. Towards this aim, FedGAT leverages synthetic MR images produced by a generative autoregressive transformer (GAT) as a privacy-preserving means of cross-site knowledge transfer. To examine the importance of the proposed GAT prior, we compared the fidelity of synthetic MR images generated by the GAT prior against those generated by adversarial (GAN) or diffusion (DDPM) priors. Fidelity of multi-site MR images generated by different priors as measured by FID are displayed in Fig. 5, and sample synthetic MR images are shown in Fig. 6. We find that GAT achieves the highest image fidelity, as manifested by its consistently lower FID scores compared to variants, and by its improved visual realism apparent in sample images. While GAN shows structural artifacts and suboptimal contrast and DDPM occasionally shows elevated noise levels, GAT achieves high image quality with lower artifacts and noise. We then assessed whether this improvement in synthetic image fidelity translates onto benefits in reconstruction performance. For this purpose, performance of reconstruction models trained on synthetic images from the GAT, GAN and DDPM priors were compared as listed in Table 5, in reference to a benchmark model that was trained using actual multi-site MR images. We find that the GAT prior yields significantly higher reconstruction performance than variants based on GAN and DDPM priors (p$<$0.05), and it attains the most competitive performance against the benchmark model. These results corroborate that the superior image fidelity of the proposed GAT prior translates onto improving the training of site-specific reconstruction models.
Next, we examined the efficacy of the knowledge-transfer approach in FedGAT that transfers information on the distribution of multi-site data via synthetic MR images. For this purpose, FedGAT was compared against a ‘w model agg.’ variant that simultaneously transferred knowledge and built reconstruction models via conventional model aggregation [36]. The comparisons were performed in model-homogeneous settings to permit use of model aggregation, and to avoid interactions between model architecture and data distribution across sites that could bias assessments. Performance metrics are listed in Table 6. FedGAT yields superior performance against ‘w model agg.’ in all cases ((p$<$0.05), except for within-site reconstructions at Site 3 where ‘w model agg.’ yields slightly higher SSIM. These results indicate that the decoupling of knowledge transfer and model building stages in FedGAT offers performance benefits against a coupled model aggregation procedure.
6 Discussion
Here we introduced a model-agnostic FL technique for MRI reconstruction that supports model-heterogeneous settings by decoupling cross-site knowledge transfer from building of reconstruction models. In particular, FedGAT mediates knowledge transfer among sites by sharing a global GAT prior that uses autoregressive predictions across multiple spatial scales to generate synthetic MR images, and a site prompt for precise control over the synthetic image distribution. Previous FL studies aiming to build generative priors have predominantly adopted adversarial approaches that can suffer from poor image quality due to training instabilities [20, 18, 74], or diffusion approaches that can suffer from extensive inference times and residual image noise [75, 76, 77, 78]. In comparison, the autoregressive nature of GAT offers improved image fidelity over adversarial and diffusion priors as indicated by our results. To curate a hybrid multi-site MRI dataset, the shared GAT prior is employed at each site to efficiently generate high-fidelity synthetic MRI datasets from remaining sites. Afterwards, site-specific reconstruction models are locally trained on this hybrid dataset to improve generalization while enjoying total freedom in architecture selection. Therefore, the proposed method holds great promise for advancing the practicality and scalability of FL in multi-institutional MRI studies.
Figure 5: Fidelity of synthetic MR images generated via GAN, DDPM and GAT priors. Bar plots of Frechet inception distance (FID) show mean±$\pm$±std across the test set.
Figure 6: Samples of synthetic MR images from GAN, DDPM and GAT priors, along with samples of local MR images at separate sites.
Table 5: Reconstruction performance of models trained on synthetic MR images generated by GAN, DDPM, and GAT priors, in reference to a benchmark model trained on actual MR images. Boldface indicates the top-performing generative prior in each task.
Within-site Site 1 (BraTS) Site 2 (fastMRI) Site 3 (MIDAS)
reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
Actual 4x 47.38$\pm$2.61 99.61$\pm$0.15 40.00$\pm$3.12 96.76$\pm$1.36 34.29$\pm$2.63 93.95$\pm$2.06
data 8x 40.89$\pm$2.78 98.08$\pm$0.85 35.02$\pm$3.00 94.14$\pm$2.05 29.74$\pm$2.63 88.78$\pm$3.36
GAN 4x 44.13$\pm$3.08 98.86$\pm$0.49 38.68$\pm$3.39 96.20$\pm$1.68 33.06$\pm$2.62 90.65$\pm$3.83
8x 39.59$\pm$2.70 95.44$\pm$1.75 34.29$\pm$3.06 93.08$\pm$2.53 28.38$\pm$2.53 85.66$\pm$4.55
DDPM 4x 45.91$\pm$2.84 99.46$\pm$0.21 38.60$\pm$3.36 96.41$\pm$1.47 33.74$\pm$2.48 92.48$\pm$2.56
8x 39.44$\pm$2.87 96.36$\pm$1.46 34.07$\pm$2.78 93.44$\pm$2.15 28.44$\pm$2.45 86.60$\pm$3.40
GAT 4x 47.17$\pm$2.56 99.60$\pm$0.13 39.81$\pm$3.41 96.47$\pm$1.64 33.80$\pm$2.62 93.35$\pm$2.45
8x 41.30$\pm$2.89 98.59$\pm$0.59 35.09$\pm$3.01 94.22$\pm$2.22 29.14$\pm$2.59 88.85$\pm$3.45
Across-site Site 1 (BraTS) Site 2 (fastMRI) Site 3 (MIDAS)
reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
Actual 4x 37.14$\pm$4.06 95.35$\pm$2.24 40.83$\pm$7.05 96.78$\pm$3.19 43.69$\pm$4.68 98.18$\pm$1.73
data 8x 32.38$\pm$3.87 91.46$\pm$3.86 35.31$\pm$6.20 93.43$\pm$5.26 37.96$\pm$4.12 96.11$\pm$2.52
GAN 4x 35.87$\pm$4.13 93.43$\pm$4.05 38.59$\pm$6.23 94.76$\pm$4.93 41.41$\pm$4.23 97.53$\pm$1.81
8x 31.33$\pm$4.08 89.37$\pm$5.23 33.98$\pm$6.18 90.55$\pm$5.98 36.94$\pm$3.92 94.26$\pm$2.47
DDPM 4x 36.17$\pm$3.83 94.48$\pm$2.89 39.82$\pm$6.64 95.97$\pm$3.93 42.25$\pm$4.80 97.97$\pm$1.82
8x 31.25$\pm$3.85 90.02$\pm$4.45 33.94$\pm$6.12 91.48$\pm$5.54 36.75$\pm$3.90 94.90$\pm$2.35
GAT 4x 36.80$\pm$4.28 94.91$\pm$2.61 40.48$\pm$7.17 96.48$\pm$3.58 43.49$\pm$4.76 98.04$\pm$1.95
8x 32.11$\pm$4.09 91.53$\pm$3.95 35.22$\pm$6.67 93.72$\pm$5.46 38.20$\pm$4.29 96.40$\pm$2.72
Table 6: Ablation study on the efficacy of the knowledge transfer approach in FedGAT.
Within-site Site 1 (BraTS) Site 2 (fastMRI) Site 3 (MIDAS)
reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
FedGAT 4x 47.63$\pm$2.65 99.64$\pm$0.14 39.73$\pm$3.31 96.46$\pm$1.56 33.59$\pm$2.63 93.53$\pm$2.10
8x 41.60$\pm$2.85 98.78$\pm$0.54 34.58$\pm$3.36 94.14$\pm$2.33 28.85$\pm$2.62 87.75$\pm$3.34
w model agg. 4x 46.65$\pm$2.86 99.58$\pm$0.14 38.52$\pm$3.62 95.50$\pm$2.32 33.14$\pm$2.52 92.80$\pm$2.75
8x 40.80$\pm$3.12 97.96$\pm$0.82 32.94$\pm$3.52 91.83$\pm$3.59 28.77$\pm$2.23 87.85$\pm$3.43
Across-site Site 1 (BraTS) Site 2 (fastMRI) Site 3 (MIDAS)
reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
FedGAT 4x 36.71$\pm$3.01 95.02$\pm$1.96 40.28$\pm$2.65 96.39$\pm$1.81 43.52$\pm$2.83 98.06$\pm$1.04
8x 32.02$\pm$2.95 91.53$\pm$3.07 35.52$\pm$2.84 93.68$\pm$2.55 37.98$\pm$2.97 95.65$\pm$1.94
w model agg. 4x 35.83$\pm$3.12 94.15$\pm$2.54 39.89$\pm$2.70 96.19$\pm$1.95 42.59$\pm$3.26 97.54$\pm$1.64
8x 30.85$\pm$2.95 89.84$\pm$3.51 34.78$\pm$2.71 92.91$\pm$2.49 36.87$\pm$3.33 94.89$\pm$2.60
The distribution of MRI datasets collected in separate sites inevitably reflects the variations in imaging protocols or scanners across sites [79]. When straightforward averaging of model weights is used for aggregation, this distributional variability can compromise the sensitivity of the global model to site-specific image attributes in the conventional FL framework. Recent FL methods on MRI reconstruction have aimed at addressing this limitation via personalization strategies such as latent-space alignment [30], partial network aggregation [32], test-time adaptation [31], or feature map normalization [42]. While these strategies allow a modest degree of model attuning to individual sites, a substantial portion of the global model must still be shared across sites to ensure adequate knowledge transfer. In contrast to existing FL methods that share reconstruction model weights, FedGAT transfers knowledge among sites by building a global generative prior that can later generate synthetic MR images from desired sites. This stark difference enables FedGAT to operate seamlessly under demanding model-heterogeneous settings, where no component of reconstruction models are shared among sites. In turn, FedGAT can sustain a higher-degree of site specialization in reconstruction models than permitted by previous approaches.
Devised to accurately capture an underlying distribution, generative priors synthesize MR images depicting random anatomical structures that do not correspond to the precise anatomy of any actual subject [75, 80]. In FedGAT, such synthetic MR images from separate sites are employed in training of site-specific reconstruction models to improve generalization across sites. Our assessments based on FID scores indicate that the synthetic MR images generated by the GAT prior attain a level of quality and realism closely mimicking actual MR images. Furthermore, reconstruction models trained on these synthetic MR images yield high performance when independently evaluated on actual MRI acquisitions in the test sets. Still, in the absence of anatomical guidance, the stochastic nature of generative priors naturally induce a degree of susceptibility to artificial features in synthetic MR images [81]. Our observations indicate that the downstream reconstruction models trained on synthetic MR images in this study are not unduly influenced by undesirable hallucinations. When necessary, hallucinations might be alleviated by building reconstruction models equipped with physics-driven modules and learning objectives [82], or building generative priors via architectures that offer enhanced capture of contextual relationships among tissue signals [83, 84]. It remains important future work to validate the proposed method and its anatomical fidelity on broader patient cohorts.
FL frameworks help mitigate patient privacy risks by exchanging model weights rather than actual MRI data across sites. Regardless, security concerns can still be present in the presence of adversaries that attempt to corrupt model updates, hence compromising the accuracy of reconstructed images [25]. In addition, third-party inference attacks on trained reconstruction models might be able to leak sensitive information regarding the set of MR images used in training sets by probing the model weights [25]. Previous methods based on the conventional FL framework directly communicate weights of a global reconstruction model across sites, so they are more susceptible to compromises in reconstruction fidelity due to training and inference attacks. In contrast, FedGAT privately builds reconstruction models at individual sites without ever communicating their weights, so it can show a degree of improved immunity against such attacks. Potential concerns related to the global GAT prior might be alleviated by incorporating differential privacy objectives in training of generative priors [85, 86]. Future research is warranted to systematically evaluate the privacy-preserving capabilities of FedGAT.
7 Conclusion
Here, we introduced a novel model-agnostic federated learning technique that enables multiple sites with distinct architectural preferences to collaborate in building MRI reconstruction models. To support model-heterogeneous settings, FedGAT decouples a first tier to learn a global generative prior that captures the distribution of multi-site MR images, from a second tier to build site-specific albeit generalizable reconstruction models with the aid of synthetic datasets locally generated by the global prior. Our experiments clearly demonstrate that FedGAT achieves superior performance compared to state-of-the-art FL baselines in both within-site and across-site reconstructions. These findings indicate that FedGAT promotes effective knowledge transfer across sites while remaining agnostic to architectures of reconstruction models, hence overcoming a critical barrier in supporting flexible collaborations across diverse institutions.

### v2

@@
 Although learning-based models hold great promise for MRI reconstruction, single-site models built on limited local datasets often suffer from poor generalization.
-This challenge has spurred interest in collaborative model training on multi-site datasets via federated learning (FL)– a privacy-preserving framework that aggregates model updates instead of sharing imaging data.
-Conventional FL builds a global model by aggregating locally trained model weights, inherently constraining all sites to a homogeneous model architecture.
-This rigid homogeneity requirement forces sites to forgo architectures tailored to their compute infrastructure and application-specific demands.
-Consequently, existing FL methods for MRI reconstruction fail to support model-heterogeneous settings, where individual sites are allowed to use distinct architectures.
-To overcome this fundamental limitation, here we introduce FedGAT, a novel model-agnostic FL technique based on generative autoregressive transformers.
-FedGAT decentralizes the training of a global generative prior that captures the distribution of multi-site MR images.
-For enhanced fidelity, we propose a novel site-prompted GAT prior that controllably synthesizes MR images from desired sites via autoregressive prediction across spatial scales.
-Each site then trains its site-specific reconstruction model–using its preferred architecture–on a hybrid dataset comprising the local MRI dataset and GAT-generated synthetic MRI datasets for other sites.
-Comprehensive experiments on multi-institutional datasets demonstrate that FedGAT supports flexible collaborations while enjoying superior within-site and across-site reconstruction performance compared to state-of-the-art FL baselines.
+This challenge has spurred interest in collaborative model training on multi-site datasets via federated learning (FL)—a privacy-preserving framework that aggregates model updates instead of sharing imaging data.
+Conventional FL aggregates locally trained model weights into a global model, inherently constraining all sites to use a homogeneous model architecture.
+This rigidity forces sites to compromise on architectures tailored to their compute resources and application-specific needs, making conventional FL unsuitable for model-heterogeneous settings where each site may prefer a distinct architecture.
+To overcome this limitation, we introduce FedGAT, a novel model-agnostic FL technique based on generative autoregressive transformers.
+FedGAT decentralizes the training of a global generative prior that learns the distribution of multi-site MR images.
+For high-fidelity synthesis, we propose a novel site-prompted GAT prior that controllably synthesizes realistic MR images from desired sites via autoregressive prediction across spatial scales.
+Each site then trains its own reconstruction model–using an architecture of its choice–on a hybrid dataset augmenting its local MRI dataset with GAT-generated synthetic MR images emulating datasets from other sites.
+This hybrid training strategy enables site-specific reconstruction models to generalize more effectively across diverse data distributions while preserving data privacy.
+Comprehensive experiments on multi-institutional datasets demonstrate that FedGAT enables flexible, model-heterogeneous collaborations and achieves superior within-site and cross-site reconstruction performance compared to state-of-the-art FL baselines.
 {IEEEkeywords}
-MRI, reconstruction, federated learning, model agnostic, generative, autoregressive, transformer.
+MRI reconstruction, federated learning, generative, autoregressive transformer.
 […]
@@
 Yet, characteristically long scan times in MRI exams can disrupt clinical workflows and reduce patient comfort [1].
-A mainstream approach to improve scan efficiency relies on reconstruction of MR images from undersampled acquisitions [2, 3].
+A mainstream approach to improve scan efficiency relies on reconstruction of MR images from undersampled acquisitions [2].
 Learning-based reconstruction models have shown particular promise in recovering high-quality images from undersampled k-space data [3, 4, 5, 6, 7, 8, 9, 10, 11].
@@
 Federated learning (FL) is an emergent paradigm for collaborative model training in privacy-sensitive domains such as medical imaging [26, 27].
-By exchanging model weights as opposed to raw imaging data, FL enables knowledge transfer among sites while alleviating potential privacy risks [29].
+By exchanging model weights as opposed to raw imaging data, FL enables knowledge transfer among sites while alleviating potential privacy risks [28, 29].
 Recognizing the transformative potential of FL, several recent studies on MRI reconstruction have proposed FL methods for building models on multi-site datasets [30, 31, 32, 33, 34, 35].
@@
 The conventional FL framework promises enhanced generalization in MRI reconstruction, while imposing a strict requirement for all participating sites to adopt a homogeneous model architecture as the basis of model aggregation [37].
-However, imaging sites often differ substantially in terms of their computational resources and the complexity of their reconstruction tasks as influenced by variations in distribution of local datasets.
+However, imaging sites often differ substantially in terms of their computational resources and the characteristics of their reconstruction tasks as influenced by variations in the distribution of local datasets.
 These differences frequently lead sites to favor distinct model architectures, ranging from convolutional networks that can be preferred to attain high local precision under relatively low compute budgets [30, 31], transformers that can be preferred to boost sensitivity to long-range contextual features when high compute budgets are available [38, 39], to physics-driven unrolled networks that can be preferred to boost model reliability when training sets are relatively scarce albeit moderate-to-high compute budgets are available [34, 33].
-Consequently, the architectural constraints in conventional FL refrain sites from leveraging models optimized for their specific requirements.
+Consequently, the architectural constraints in conventional FL refrain sites from leveraging models optimized for their specific needs.
 This apparent lack of architectural freedom has important practical implications including discouraging resource-constrained sites from participating in collaborations, compromising performance at sites with more advanced resources due to use of relatively simple models, and ultimately stifling innovations in MRI reconstruction models.
@@
 In this study, we introduce a novel FL technique based on generative autoregressive transformers (FedGAT), which to our knowledge is the first model-agnostic FL approach in the literature for MRI reconstruction.
-Unlike conventional FL methods, FedGAT supports model-heterogeneous settings by decoupling the process of cross-site knowledge transfer on data distributions, from the process of building reconstruction models (Fig.
+Unlike conventional FL methods that require consistent model architectures across sites, FedGAT supports model-heterogeneous settings via a two-tier strategy that decouples cross-site knowledge transfer, from building of reconstruction models (Fig.
 1).
-It mediates knowledge transfer via a global generative prior designed to synthesize multi-site MR images.
-To avoid the limitations of existing adversarial priors (i.e., training instability and poor quality) and diffusion priors (i.e., suboptimal convergence and heavy compute burden), here we propose a novel GAT prior that autoregressively synthesizes MR images by predicting feature maps across spatial scales, guided by a site prompt to preserve site-specific characteristics.
-Afterwards, site-specific reconstruction models are locally trained on a hybrid dataset, combining local MRI datasets available at each site with synthetic datasets from remaining sites generated by the GAT prior.
-Comprehensive demonstrations on multi-institutional MRI datasets indicate that FedGAT functions effectively in model-heterogeneous settings, overcoming a significant challenge in enabling privacy-preserving collaborations across diverse institutions.
+To mediate knowledge transfer without sharing private data, in the first tier, FedGAT trains a global generative prior —a generative autoregressive transformer (GAT)— capable of synthesizing realistic multi-site MR images.
+Although adversarial and diffusive methods have been common in medical image synthesis [41, 42, 43], adversarial priors are amenable to training instabilities [44], diffusion priors are amenable to high computational complexity [45], potentially compromising convergence, image fidelity and practical utility in FL settings [31, 46].
+To address these limitations, we propose a novel GAT prior that autoregressively predicts feature maps at multiple spatial scales, conditioned on site-specific prompts to preserve the unique characteristics of each site’s data.
+The proposed prior balances efficiency and fidelity while capturing diverse data distributions across sites within a unified model.
+Having a trained copy of the GAT prior, in the second tier, each site builds its own reconstruction model using a hybrid dataset: augmenting its own local MRI dataset with synthetic images generated by the GAT prior to emulate datasets from other sites.
+This hybrid training ensures that site-specific reconstruction models are exposed to broader data variability while preserving data privacy.
+Comprehensive experiments on multi-institutional MRI datasets demonstrate that FedGAT enables effective federated learning in model-heterogeneous settings, representing an important advance toward privacy-preserving collaborations across diverse institutions.
 Code for FedGAT is available at https://github.com/icon-lab/FedGAT.
@@
   • To strike a balance between generalization and site-specific performance, FedGAT leverages a training procedure where reconstruction models pre-trained on local datasets are fine-tuned on a hybrid dataset including synthetic datasets from other sites.
-Figure 1: FedGAT devises a two-tier strategy to collaboratively train heterogeneous models for MRI reconstruction.
-(a) The first tier conducts decentralized training of a global prior $\text{GAT}_{\theta_{\text{GAT}}}$ that captures the distribution of multi-site MR images.
-The global prior is built on a novel generative autoregressive transformer that formulates image generation as an autoregressive process across growing spatial scales, and that employs a site prompt sp for controllable image synthesis from desired sites.
-(b) The second tier conducts local training of site-specific reconstruction models $H^{k}_{{\phi}^{k}}$ ($k$: site index).
-To achieve an optimal balance between generalization and site-specific performance, training at each site is performed on a hybrid dataset comprising the local dataset and synthetic datasets from other sites generated via the GAT prior.
+Figure 1: FedGAT implements a two-tier strategy to collaboratively train heterogeneous models for MRI reconstruction across multiple imaging sites.
+(a) The first tier conducts decentralized training of a global prior $\text{GAT}_{\theta_{\text{GAT}}}$ that learns the distribution of multi-site MR images.
+This global prior is based on a novel generative autoregressive transformer that formulates image generation as an autoregressive process across growing spatial scales, guided by a site prompt sp to preserve site-specific characteristics.
+(b) The second tier performs local training of site-specific reconstruction models $H^{k}_{{\phi}^{k}}$ ($k$: site index).
+Each model is trained on a hybrid dataset that combines the local MRI dataset with GAT-generated synthetic images representing datasets from other sites, enabling broader generalization while maintaining data privacy.
 2 Related Work
@@
 Learning-based models promise improved performance and reliability over traditional methods in MRI reconstruction [4].
-Yet, realizing this promise requires copious amounts of training data including diverse samples, since learning-based models poorly represent rare features and pathologies in their training sets [40].
+Yet, realizing this promise requires copious amounts of training data including diverse samples, since learning-based models poorly represent rare features and pathologies in their training sets [47, 48, 49].
 Unfortunately, curating broad training sets is unfeasible in many applications given the high costs of scanning and subject recruitment [50].
@@
 Several recent MRI studies have successfully adopted this conventional FL framework to boost generalization performance of reconstruction models [30, 31, 32].
-Further improvements in site-specific performance have also been sought by employing personalized FL strategies to cope with differences in the distribution of MRI datasets across sites, such as partial model aggregation [32], test-time adaptation [31], and feature map normalization [42].
+Further improvements in site-specific performance have also been sought by employing personalized FL strategies to cope with differences in the distribution of MRI datasets across sites, including partial model aggregation [32], test-time adaptation [31], and feature map normalization [51].
 Despite their apparent benefits, previous FL methods for MRI reconstruction constrain all participating sites to adopt a homogeneous model architecture, which serves as the foundation for aggregating model weights [37].
@@
 At each site, a site-specific reconstruction model is then locally trained on a hybrid dataset comprising both the local dataset and synthetic datasets from other sites generated via the global prior.
-These technical advances enable FedGAT to operate seamlessly in model-heterogeneous settings, overcoming a critical barrier towards privacy-preserving collaborations across diverse institutions.
+These technical advances enable FedGAT to operate effectively in model-heterogeneous settings, overcoming a critical barrier towards collaborations across diverse institutions.
 3 Theory
@@
 As such, reconstruction for a test subject at site $k$ is formulated as:
-$\widehat{\mathbf{x}}^{k}=H_{\phi^{*}}(\mathcal{F}^{-1}\mathbf{y}^{k},\mathcal{A}^{k}),$ (5)
-where $\widehat{\mathbf{x}}^{k}$ is the reconstructed image, and $\mathbf{y}^{k}$ are the respective undersampled k-space data.
+$\widehat{\mathbf{x}}\hskip 1.00006pt^{k}=H_{\phi^{*}}(\mathcal{F}^{-1}\mathbf{y}^{k},\mathcal{A}^{k}),$ (5)
+where $\widehat{\mathbf{x}}\hskip 1.00006pt^{k}$ is the reconstructed image, and $\mathbf{y}^{k}$ are the respective undersampled k-space data.
 3.2 Federated Generative Autoregressive Transformers
 FedGAT is a model-agnostic FL technique that enables collaborative training of heterogeneous model architectures across sites by leveraging a decoupled two-tier strategy.
-During the first tier, using a server-client topology, a site-prompted global prior based on generative autoregressive transformers (GAT) is decentrally trained so as to capture the distribution of multi-site MR images (Fig.
+During the first tier, using a server-client topology, a site-prompted global prior based on generative autoregressive transformers (GAT) is trained so as to capture the distribution of multi-site MR images (Fig.
 1a).
@@
 This module expresses generation of multi-scale discrete token maps as an autoregressive process across spatial scales, wherein a higher-scale token map $\mathbf{f}_{s}$ is predicted given the collection of token maps at earlier scales $\mathbf{f}_{<s}$:=assign\,:=\,:=$\{\mathbf{f}_{1},...,\mathbf{f}_{s-1}\}$.
-To preserve site-specific attributes in synthetic MR images, a site-token $\mathbf{st}$ derived from a site prompt is included in the autoregressive transformer module.
+To preserve site-specific attributes in synthetic MR images, a site-prompt $\mathbf{sp}(k)$ is learned via a gated MLP (gMLP) network from a one-hot encoding of site index, and used to initialize the site token $\mathbf{st}$ in the autoregressive transformer module.
 VAE Decoder Module: The VAE decoder aims to map discrete multi-scale token maps $\{\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{S}\}$ back on the corresponding MR image $\mathbf{x}$ from which they were derived.
@@
 To obtain a correlate for the joint distribution of $\mathbf{f}_{1},..,\mathbf{f}_{S}$, here we propose to build an autoregressive prior on multi-scale discrete token maps via the autoregressive transformer module.
-In particular, the transformer module aims to predict the sequence of tokens $\mathbf{f}_{s}$ at the $s$th scale, given as input the sequence of tokens aggregated across lower scales $\mathbf{f}_{<s}$:=assign\,:=\,:=$\{\mathbf{f}_{1},...,\mathbf{f}_{s-1}\}\in[V]^{T_{s-1}}$ ($T_{s-1}=\sum_{o=1}^{s-1}h_{o}w_{o}$ denoting the number of tokens at lower scales), and a learnable site-prompt $\mathbf{sp}(k)\in\mathbb{R}^{1\times d}$ used to initialize a site token, $\mathbf{st}$=\,=\,=$\mathbf{sp}(k)$.
-To maintain spatial correspondence between the token maps at consecutive scales, the token maps in $\mathbf{f}_{<s}$ are individually embedded after upsampling to the dimensions of the subsequent scale:
+In particular, the transformer module aims to predict the sequence of tokens $\mathbf{f}_{s}$ at the $s$th scale, given as input the sequence of tokens aggregated across lower scales $\mathbf{f}_{<s}$:=assign\,:=\,:=$\{\mathbf{f}_{1},...,\mathbf{f}_{s-1}\}\in[V]^{T_{s-1}}$ ($T_{s-1}=\sum_{o=1}^{s-1}h_{o}w_{o}$ denoting the number of tokens at lower scales).
+To maintain spatial correspondence across consecutive scales, the token maps in $\mathbf{f}_{<s}$ are individually embedded after upsampling to the dimensions of the subsequent scale:
 $\mathbf{e}_{o}=\mathrm{Lin}(\mathrm{Up}_{o+1}(\mathbf{f}_{o})),\mbox{ for }o\in[1\mbox{ }s-1],$ (11)
 where $\mathbf{e}_{o}\in\mathbb{R}^{h_{o+1}w_{o+1}\times d}$ and $\mathbf{e}_{<s}$:=assign\,:=\,:=$\{\mathbf{e}_{1},...,\mathbf{e}_{s-1}\}\in\mathbb{R}^{(T_{s}-1)\times d}$.
-Afterwards, the site token is pooled with a learnable projection of these embedded token maps, and additively combined with learnable position and scale encodings:
+To enable the transformer module to synthesize feature maps aligned with site-specific data distributions, a learnable site-prompt $\mathbf{sp}(k)\in\mathbb{R}^{1\times d}$ is generated by projecting a one-hot encoding of site index ($\mathbb{1}_{K}(k)$) through a gated multi-layer perceptron (gMLP) network:
+$\mathbf{g}$ $=\mathrm{GELU}(\mathrm{Lin}(\mathbb{1}_{K}(k))),$ (12)
+$\mathbf{g}_{u},\mathbf{g}_{v}$ $=\mathrm{Split}(\mathbf{g}),$ (13)
+$\mathbf{sp}(k)$ $=\mathrm{Lin}(\mathrm{Lin}(\mathrm{LN}(\mathbf{g}_{u}))\odot\mathbf{g}_{v}),$ (14)
+where GELU is Gaussian error linear unit activation, $\mathbf{g}$ is the hidden activation, $\mathbf{g}_{u}$ and $\mathbf{g}_{v}$ are the channel-split components of $\mathbf{g}$, Lin denotes a linear layer, and LN denotes layer normalization.
+The site prompt is then used to initialize a site token as $\mathbf{st}$=\,=\,=$\mathbf{sp}(k)$.
+Afterwards, the site token is pooled with a learnable projection of the embedded token maps and additively combined with learnable position and scale encodings:
 $\mathbf{h}_{0}=\Big{[}\big{[}\mathbf{st};\,\,\mathbf{e}_{<s}\cdot\mathbf{E}_{\text{tok}}\big{]}+\mathbf{E}_{\text{pos}}+\mathbf{E}_{\text{sca}};\,\,\mathbf{m}\Big{]},$ (15)
@@
 In Eq.
-12, $\mathbf{m}$=\,=\,=$0^{(T_{S}-T_{s})\times d}$ is a zero matrix that masks tokens from spatial scales succeeding the $(s-1)$th scale to ensure causal progression from lower to higher scales.
-To predict $\mathbf{f}_{s}$, the hidden representation $\mathbf{h}_{0}$ is projected through $L$ transformer blocks, each containing multi-head self-attention (MHSA) and multi-later perceptron (MLP) layers interleaved with normalization and residual connections [55].
+15, $\mathbf{m}$=\,=\,=$0^{(T_{S}-T_{s})\times d}$ is a zero matrix that masks tokens from spatial scales succeeding the $(s-1)$th scale to ensure causal progression from lower to higher scales.
+To predict $\mathbf{f}_{s}$, the hidden representation $\mathbf{h}_{0}$ is projected through $L$ transformer blocks, each containing multi-head self-attention (MHSA) and MLP layers interleaved with normalization and residual connections [64].
 Yet, unlike normalization layers in conventional transformers, here we employ adaptive layer normalization (AdaLN) in order to induce site-specific processing of hidden representations [70].
@@
 The transformer module generates a random set of multi-scale discrete token maps, sequentially across growing spatial scales based on the autoregressive mapping in Eq.
-20.
+23.
 To introduce stochasticity during inference, however, nucleus sampling [71] is employed instead of the maximum-probability sampling procedure in Eqs.
-18-19.
+21-22.
 Accordingly, the logit values across $V$ codebook vectors are ordered separately for each token in $\hat{\mathbf{f}}_{s}$, the top-5 percentile of values are subjected to softmax to obtain a probability mass function, and the predicted token is randomly drawn from this function:
@@
 The procedural details and loss functions used in each tier are discussed below.
-Training of the GAT prior: FedGAT decentrally trains a global, site-prompted GAT prior that captures the distribution of multi-site MR images, which is later used to generate synthetic MR images from desired sites.
+Tier 1 - Training of the GAT prior: FedGAT decentrally trains a global, site-prompted GAT prior that captures the distribution of multi-site MR images, which is later used to generate synthetic MR images from desired sites.
 The proposed GAT prior embodies VAE and autoregressive transformer modules as described in Section 3.2.1.
@@
 Algorithm 2 Training of reconstruction models
-Training of reconstruction models: Following the first tier, each individual site trains a site-specific reconstruction model $H^{k}_{\phi^{k}}$ of preferred architecture, with the aid of the global GAT prior (Alg.
+Tier 2 - Training of reconstruction models: Following the first tier, each individual site trains a site-specific reconstruction model $H^{k}_{\phi^{k}}$ of preferred architecture, with the aid of the global GAT prior (Alg.
 2).
@@
 FedGAT was comparatively demonstrated against non-FL benchmarks (Central, Single) along with distillation-based (FedMD, FedDF) and generative (FedGIMP, FedDDA) FL baselines.
+Since previous FL methods for MRI reconstruction were not specifically designed to handle model-heterogeneous settings, we either adopted model-agnostic baselines from non-reconstruction tasks (FedMF, FedDG, FedDDA), or modified reconstruction baselines to accommodate model heterogeneity when feasible (FedGIMP).
 Central: A non-FL benchmark for across-site reconstruction performance was considered that pools MRI datasets across sites to centrally train a global reconstruction model [24].
@@
 FedDF: A distillation-based FL baseline was considered that supports model-heterogeneous settings via an auxiliary public dataset [75].
-To adopt FedDF for MRI reconstruction, distillation between site-specific reconstruction models and a global model were performed based on predictions of reconstructed images, instead of class logits as in the original implementation for classification tasks [66].
-At individual sites, site-specific reconstruction models were initialized via distillation from the global model (taken as MoDL-5), and then trained along with the copy of the global model using a proximal loss term [67].
+To adopt FedDF for MRI reconstruction, distillations between site-specific reconstruction models and a global model were performed based on predictions of reconstructed images, instead of class logits as in the original implementation for classification tasks [75].
+At individual sites, site-specific reconstruction models were initialized via distillation from the global model (taken as MoDL5), and then trained along with the copy of the global model using a proximal loss term [76].
 Training was performed on a 840-sample mixture of the local data with independent auxiliary data downloaded from a public repository [77].
-On the server, copies of the global model locally trained at individual sites were aggregated via distillation on a 840-sample set of auxiliary data.
+On the server, copies of the global model locally trained at individual sites were aggregated via distillation on a 840-sample set of the auxiliary data.
 FedMD: A distillation-based FL baseline was considered that supports model-heterogeneous settings via an auxiliary public dataset [78].
@@
 A site index was included in FedDDA for controllable generation of multi-site MR images [51].
-FedDDA was imeplemented using the same two-tier training strategy as in FedGAT to support model-heterogeneous settings, where site-specific reconstruction models were trained using a hybrid dataset containing both local MRI data and synthetic data from remaining sites generated by FedDDA.
-Table 1: Reconstruction performance for single-coil datasets.
-Each site employs an entirely distinct model type, with Site $k$ (dataset-type) denoting the site index, the local dataset, and the type of reconstruction model.
+FedDDA was implemented using the same two-tier training strategy as in FedGAT to support model-heterogeneous settings, where site-specific reconstruction models were trained using a hybrid dataset containing both local MRI data and synthetic data from remaining sites generated by FedDDA.
+Table 1: Reconstruction performance for coil-combined datasets (Site 1: BraTS, Site 2: fastMRI, Site 3: MIDAS).
+Site $k$ (type) denotes site index and type of reconstruction model.
 Metrics are listed as mean±$\pm$±std across the test set, for within-site (upper panel) and across-site (lower panel) reconstructions at R=4x-8x.
@@
 Benchmark methods for within-site reconstruction (Single), and for cross-site reconstruction (Central) are underlined.
-Within-site Site 1 (BraTS-MoDL) Site 2 (fastMRI-rGAN) Site 3 (MIDAS-D5C5)
-reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
-non-FL Single 4x 48.06$\pm$3.33 99.72$\pm$0.11 35.83$\pm$2.51 90.64$\pm$3.87 29.04$\pm$1.84 76.77$\pm$4.08
-8x 42.46$\pm$3.15 99.03$\pm$0.45 33.18$\pm$2.26 86.58$\pm$4.64 27.79$\pm$2.05 74.20$\pm$3.97
-Central 4x 46.12$\pm$2.55 99.46$\pm$0.20 36.03$\pm$2.51 91.23$\pm$3.63 28.89$\pm$1.89 75.11$\pm$4.07
-8x 42.16$\pm$2.72 98.65$\pm$0.58 33.18$\pm$2.53 87.40$\pm$4.44 27.34$\pm$2.17 73.87$\pm$3.98
-FL methods FedDF 4x 44.42$\pm$3.76 99.34$\pm$0.28 32.88$\pm$3.22 87.95$\pm$4.41 27.62$\pm$1.90 61.23$\pm$4.32
-8x 40.06$\pm$3.23 98.23$\pm$0.77 30.78$\pm$2.73 83.45$\pm$5.51 27.48$\pm$1.93 59.62$\pm$4.93
-FedMD 4x 44.68$\pm$3.72 99.40$\pm$0.24 34.24$\pm$3.06 88.48$\pm$4.13 28.95$\pm$1.92 62.22$\pm$5.25
-8x 40.32$\pm$3.22 98.29$\pm$0.75 32.96$\pm$2.56 86.40$\pm$4.66 27.79$\pm$2.07 60.19$\pm$5.07
-FedGIMP 4x 47.61$\pm$2.86 99.63$\pm$0.15 35.45$\pm$2.48 90.27$\pm$3.99 28.68$\pm$1.90 74.91$\pm$4.42
-8x 42.14$\pm$2.71 98.62$\pm$0.54 32.74$\pm$2.40 85.86$\pm$4.75 26.82$\pm$2.17 73.96$\pm$4.08
-FedDDA 4x 47.50$\pm$2.81 99.61$\pm$0.15 35.67$\pm$2.51 90.30$\pm$3.94 28.61$\pm$1.79 75.67$\pm$4.15
-8x 41.76$\pm$2.90 98.47$\pm$0.63 32.79$\pm$2.45 86.02$\pm$4.84 27.10$\pm$2.18 73.08$\pm$4.07
-FedGAT 4x 47.63$\pm$2.65 99.64$\pm$0.14 35.89$\pm$2.50 90.69$\pm$3.86 31.88$\pm$2.03 77.48$\pm$4.94
-8x 41.60$\pm$2.85 98.78$\pm$0.54 33.23$\pm$2.36 86.45$\pm$0.67 29.62$\pm$2.11 74.88$\pm$4.29
-Across-site Site 1 (BraTS-MoDL) Site 2 (fastMRI-rGAN) Site 3 (MIDAS-D5C5)
-reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
-non-FL Single 4x 34.45$\pm$3.57 91.44$\pm$4.05 33.02$\pm$2.30 66.49$\pm$7.26 29.67$\pm$2.28 75.67$\pm$6.93
-8x 30.48$\pm$3.20 85.48$\pm$5.37 30.08$\pm$2.20 59.14$\pm$6.78 28.38$\pm$2.26 73.64$\pm$6.55
-Central 4x 36.48$\pm$2.89 94.78$\pm$1.91 33.53$\pm$2.31 82.93$\pm$4.07 30.68$\pm$2.47 77.41$\pm$6.99
-8x 32.79$\pm$2.81 92.16$\pm$2.62 30.78$\pm$2.25 78.75$\pm$4.67 29.93$\pm$2.20 75.86$\pm$6.95
-FL methods FedDF 4x 34.40$\pm$3.12 93.12$\pm$3.00 33.14$\pm$2.61 71.59$\pm$6.74 29.23$\pm$2.38 72.76$\pm$7.12
-8x 30.91$\pm$2.87 90.33$\pm$3.36 30.89$\pm$2.22 64.17$\pm$7.14 28.37$\pm$2.63 72.45$\pm$8.43
-FedMD 4x 34.61$\pm$3.13 93.42$\pm$2.75 33.11$\pm$2.26 68.04$\pm$6.84 30.03$\pm$2.41 72.31$\pm$7.93
-8x 31.12$\pm$2.83 90.63$\pm$3.13 30.47$\pm$2.20 61.79$\pm$7.06 29.05$\pm$2.57 71.80$\pm$7.84
-FedGIMP 4x 36.42$\pm$3.17 94.32$\pm$2.49 33.05$\pm$2.29 67.50$\pm$7.01 30.30$\pm$2.33 76.19$\pm$6.56
-8x 31.88$\pm$2.87 90.16$\pm$3.78 29.78$\pm$2.16 57.80$\pm$6.39 28.31$\pm$2.16 71.75$\pm$6.10
-FedDDA 4x 36.19$\pm$3.13 94.41$\pm$2.24 33.11$\pm$2.21 67.78$\pm$7.00 28.59$\pm$2.07 78.13$\pm$5.12
-8x 31.79$\pm$2.86 90.75$\pm$3.07 30.15$\pm$2.16 59.18$\pm$6.93 27.73$\pm$2.13 77.92$\pm$5.11
-FedGAT 4x 36.71$\pm$3.01 95.02$\pm$1.96 34.50$\pm$2.36 75.69$\pm$5.76 30.77$\pm$2.35 78.62$\pm$6.41
-8x 32.02$\pm$2.95 91.53$\pm$3.07 31.47$\pm$2.49 71.84$\pm$6.47 29.72$\pm$2.38 78.08$\pm$6.86
-Table 2: Within-site and cross-site reconstruction performance for multi-coil datasets (f.knee denotes fastMRI-knee, f.brain denotes fastMRI-brain, UMRAM denotes the in-house dataset), where each site employs an entirely different model type.
-Within-site Site 1 (f.knee-MoDL) Site 2 (f.brain-rGAN) Site 3 (UMRAM-D5C5)
-reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
-non-FL Single 4x 33.78$\pm$2.27 93.37$\pm$1.52 36.97$\pm$3.07 97.32$\pm$1.17 31.81$\pm$1.73 90.10$\pm$2.57
-8x 29.39$\pm$2.06 83.90$\pm$3.08 33.08$\pm$3.39 94.54$\pm$2.26 28.77$\pm$1.71 85.29$\pm$3.64
-Central 4x 33.69$\pm$2.32 93.20$\pm$1.63 36.93$\pm$3.22 97.32$\pm$1.19 31.02$\pm$1.60 88.49$\pm$2.56
-8x 29.39$\pm$2.19 83.62$\pm$3.09 33.79$\pm$3.05 94.59$\pm$2.13 27.39$\pm$1.68 82.72$\pm$3.04
-FL methods FedDF 4x 32.79$\pm$2.19 90.40$\pm$2.45 33.95$\pm$2.88 96.30$\pm$1.39 30.34$\pm$1.55 88.99$\pm$2.53
-8x 27.70$\pm$2.33 78.94$\pm$4.57 28.80$\pm$2.82 90.60$\pm$3.14 27.05$\pm$1.72 83.94$\pm$3.40
-FedMD 4x 32.48$\pm$2.24 90.57$\pm$2.21 34.07$\pm$2.83 96.36$\pm$1.36 30.10$\pm$1.61 88.73$\pm$2.52
-8x 26.99$\pm$2.47 78.49$\pm$4.51 29.08$\pm$3.15 90.16$\pm$3.23 27.13$\pm$1.64 84.18$\pm$3.31
-FedGIMP 4x 33.78$\pm$2.27 93.41$\pm$1.48 31.72$\pm$3.85 94.25$\pm$4.32 30.11$\pm$2.08 88.56$\pm$2.43
-8x 28.97$\pm$2.12 83.68$\pm$3.07 28.78$\pm$3.10 90.71$\pm$5.21 26.26$\pm$2.68 82.69$\pm$3.52
-FedDDA 4x 32.53$\pm$2.33 91.75$\pm$1.82 32.69$\pm$3.46 94.71$\pm$3.44 24.47$\pm$2.57 81.20$\pm$3.47
-8x 28.09$\pm$2.22 81.90$\pm$2.93 29.63$\pm$3.28 90.85$\pm$5.46 18.38$\pm$1.97 68.73$\pm$5.73
-FedGAT 4x 34.10$\pm$2.24 97.32$\pm$1.67 35.86$\pm$2.78 96.85$\pm$1.19 31.94$\pm$1.70 90.21$\pm$2.51
-8x 29.49$\pm$2.21 83.73$\pm$3.07 31.48$\pm$2.34 92.89$\pm$2.38 28.28$\pm$1.89 85.20$\pm$3.58
-Across-site Site 1 (f.knee-MoDL) Site 2 (f.brain-rGAN) Site 3 (UMRAM-D5C5)
-reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
-non-FL Single 4x 38.40$\pm$3.26 97.32$\pm$1.67 31.17$\pm$3.22 93.03$\pm$2.57 26.18$\pm$2.62 79.42$\pm$5.73
-8x 32.39$\pm$3.04 92.63$\pm$2.94 26.82$\pm$3.15 85.13$\pm$5.06 22.64$\pm$2.30 70.31$\pm$6.85
-Central 4x 39.70$\pm$3.36 97.62$\pm$1.70 37.41$\pm$2.13 96.27$\pm$1.29 27.52$\pm$2.49 81.63$\pm$5.35
-8x 34.33$\pm$3.12 93.86$\pm$3.04 33.98$\pm$2.11 91.89$\pm$2.82 25.09$\pm$2.47 74.38$\pm$6.46
-FL methods FedDF 4x 38.03$\pm$2.99 96.54$\pm$1.41 32.30$\pm$2.93 92.69$\pm$1.76 26.68$\pm$2.55 81.09$\pm$5.31
-8x 32.05$\pm$2.80 92.25$\pm$3.12 27.68$\pm$2.80 85.12$\pm$4.50 24.04$\pm$2.48 70.60$\pm$6.94
-FedMD 4x 37.77$\pm$3.07 96.53$\pm$1.59 32.62$\pm$2.94 92.69$\pm$1.69 26.39$\pm$2.50 80.99$\pm$5.26
-8x 31.35$\pm$2.91 91.22$\pm$3.24 28.03$\pm$2.50 84.44$\pm$4.68 23.99$\pm$2.43 70.13$\pm$6.61
-FedGIMP 4x 37.87$\pm$3.13 97.23$\pm$1.50 26.57$\pm$3.13 85.71$\pm$5.69 26.00$\pm$2.49 80.03$\pm$5.68
-8x 32.01$\pm$2.98 92.52$\pm$2.91 22.85$\pm$2.93 75.96$\pm$8.28 22.20$\pm$2.29 67.89$\pm$8.82
-FedDDA 4x 35.54$\pm$3.34 94.49$\pm$2.44 27.04$\pm$3.51 84.62$\pm$4.68 22.34$\pm$2.33 72.91$\pm$6.55
-8x 30.15$\pm$3.57 88.74$\pm$3.98 23.98$\pm$3.40 75.66$\pm$5.86 19.21$\pm$2.32 60.73$\pm$9.18
-FedGAT 4x 39.25$\pm$3.09 97.50$\pm$1.52 33.95$\pm$2.34 93.45$\pm$1.95 27.17$\pm$2.58 81.24$\pm$5.63
-8x 33.52$\pm$2.98 92.90$\pm$3.11 29.20$\pm$2.28 86.58$\pm$4.23 24.45$\pm$2.55 77.63$\pm$8.77
+Within-site Site 1 (MoDL) Site 2 (rGAN) Site 3 (D5C5) Average
+[…]
+non-FL Single 4x 48.1$\pm$3.3 99.7$\pm$0.1 35.8$\pm$2.5 90.6$\pm$3.9 29.0$\pm$1.8 76.8$\pm$4.1 37.6 89.0
+8x 42.5$\pm$3.2 99.0$\pm$0.5 33.2$\pm$2.3 86.6$\pm$4.6 27.8$\pm$2.1 74.2$\pm$4.0 34.5 86.6
+Central 4x 46.1$\pm$2.6 99.5$\pm$0.2 36.0$\pm$2.5 91.2$\pm$3.6 28.9$\pm$1.9 75.1$\pm$4.1 37.0 88.6
+8x 42.2$\pm$2.7 98.7$\pm$0.6 33.2$\pm$2.5 87.4$\pm$4.4 27.3$\pm$2.2 73.9$\pm$4.0 34.2 86.7
+FL methods FedDF 4x 44.4$\pm$3.8 99.3$\pm$0.3 32.9$\pm$3.2 88.0$\pm$4.4 27.6$\pm$1.9 61.2$\pm$4.3 35.0 82.8
+8x 40.1$\pm$3.2 98.2$\pm$0.8 30.8$\pm$2.7 83.5$\pm$5.5 27.5$\pm$1.9 59.6$\pm$4.9 32.8 80.4
+FedMD 4x 44.7$\pm$3.7 99.4$\pm$0.2 34.2$\pm$3.1 88.5$\pm$4.1 29.0$\pm$1.9 62.2$\pm$5.3 36.0 83.4
+8x 40.3$\pm$3.2 98.3$\pm$0.8 33.0$\pm$2.6 86.4$\pm$4.7 27.8$\pm$2.1 60.2$\pm$5.1 33.7 81.6
+FedGIMP 4x 47.6$\pm$2.9 99.6$\pm$0.2 35.5$\pm$2.5 90.3$\pm$4.0 28.7$\pm$1.9 74.9$\pm$4.4 37.3 88.3
+8x 42.1$\pm$2.7 98.6$\pm$0.5 32.7$\pm$2.4 85.9$\pm$4.8 26.8$\pm$2.2 74.0$\pm$4.1 33.9 86.2
+FedDDA 4x 47.5$\pm$2.8 99.6$\pm$0.2 35.7$\pm$2.5 90.3$\pm$3.9 28.6$\pm$1.8 75.7$\pm$4.2 37.3 88.5
+8x 41.8$\pm$2.9 98.5$\pm$0.6 32.8$\pm$2.5 86.0$\pm$4.8 27.1$\pm$2.2 73.1$\pm$4.1 33.9 85.9
+FedGAT 4x 47.9$\pm$2.7 99.7$\pm$0.1 35.9$\pm$2.5 90.7$\pm$3.9 31.9$\pm$2.0 77.5$\pm$4.9 38.6 89.3
+8x 41.6$\pm$2.9 98.7$\pm$0.6 33.2$\pm$2.4 86.5$\pm$0.7 29.6$\pm$2.1 74.9$\pm$4.3 34.8 86.7
+Across-site Site 1 (MoDL) Site 2 (rGAN) Site 3 (D5C5) Average
+[…]
+non-FL Single 4x 34.5$\pm$3.6 91.4$\pm$4.1 33.0$\pm$2.3 66.5$\pm$7.3 29.7$\pm$2.3 75.7$\pm$6.9 32.4 77.9
+8x 30.5$\pm$3.2 85.5$\pm$5.4 30.1$\pm$2.2 59.1$\pm$6.8 28.4$\pm$2.3 73.6$\pm$6.6 29.7 72.7
+Central 4x 36.5$\pm$2.9 94.8$\pm$1.9 33.5$\pm$2.3 82.9$\pm$4.1 30.7$\pm$2.5 77.4$\pm$7.0 33.6 85.0
+8x 32.8$\pm$2.8 92.2$\pm$2.6 30.8$\pm$2.3 78.8$\pm$4.7 29.9$\pm$2.2 75.9$\pm$7.0 31.2 82.3
+FL methods FedDF 4x 34.4$\pm$3.1 93.1$\pm$3.0 33.1$\pm$2.6 71.6$\pm$6.7 29.2$\pm$2.4 72.8$\pm$7.1 32.2 79.2
+8x 30.9$\pm$2.9 90.3$\pm$3.4 30.9$\pm$2.2 64.2$\pm$7.1 28.4$\pm$2.6 72.5$\pm$8.4 30.1 75.7
+FedMD 4x 34.6$\pm$3.1 93.4$\pm$2.8 33.1$\pm$2.3 68.0$\pm$6.8 30.0$\pm$2.4 72.3$\pm$7.9 32.6 77.9
+8x 31.1$\pm$2.8 90.6$\pm$3.1 30.5$\pm$2.2 61.8$\pm$7.1 29.1$\pm$2.6 71.8$\pm$7.8 30.2 74.7
+FedGIMP 4x 36.4$\pm$3.2 94.3$\pm$2.5 33.1$\pm$2.3 67.5$\pm$7.0 30.3$\pm$2.3 76.2$\pm$6.6 33.3 79.3
+8x 31.9$\pm$2.9 90.2$\pm$3.8 29.8$\pm$2.2 57.8$\pm$6.4 28.3$\pm$2.2 71.8$\pm$6.1 30.0 73.3
+FedDDA 4x 36.2$\pm$3.1 94.4$\pm$2.2 33.1$\pm$2.2 67.8$\pm$7.0 28.6$\pm$2.1 78.1$\pm$5.1 32.6 80.1
+8x 31.8$\pm$2.9 90.8$\pm$3.1 30.2$\pm$2.2 59.2$\pm$6.9 27.7$\pm$2.1 77.9$\pm$5.1 29.9 76.0
+FedGAT 4x 36.7$\pm$4.3 95.0$\pm$2.0 34.5$\pm$2.4 75.7$\pm$5.8 30.8$\pm$2.4 78.6$\pm$6.4 34.0 83.1
+8x 32.1$\pm$4.0 91.5$\pm$3.1 31.5$\pm$2.5 71.8$\pm$6.5 29.7$\pm$2.4 78.1$\pm$6.9 31.1 80.5
+Table 2: Within-site and cross-site reconstruction performance for multi-coil datasets (Site 1: fastMRI-knee, Site 2: fastMRI-brain, Site 3: UMRAM). Each site employs a distinct model type.
+Within-site Site 1 (MoDL) Site 2 (rGAN) Site 3 (D5C5) Average
+[…]
+non-FL Single 4x 33.8$\pm$2.3 93.4$\pm$1.5 37.0$\pm$3.1 97.3$\pm$1.2 31.8$\pm$1.7 90.1$\pm$2.6 34.2 93.6
+8x 29.4$\pm$2.1 83.9$\pm$3.1 33.1$\pm$3.4 94.5$\pm$2.3 28.8$\pm$1.7 85.3$\pm$3.6 30.4 87.9
+Central 4x 33.7$\pm$2.3 93.2$\pm$1.6 36.9$\pm$3.2 97.3$\pm$1.2 31.0$\pm$1.6 88.5$\pm$2.6 33.9 93.0
+8x 29.4$\pm$2.2 83.6$\pm$3.1 33.8$\pm$3.1 94.6$\pm$2.1 27.4$\pm$1.7 82.7$\pm$3.0 30.2 87.0
+FL methods FedDF 4x 32.8$\pm$2.2 90.4$\pm$2.5 34.0$\pm$2.9 96.3$\pm$1.4 30.3$\pm$1.6 89.0$\pm$2.5 32.4 91.9
+8x 27.7$\pm$2.3 78.9$\pm$4.6 28.8$\pm$2.8 90.6$\pm$3.1 27.1$\pm$1.7 83.9$\pm$3.4 27.9 84.5
+FedMD 4x 32.5$\pm$2.2 90.6$\pm$2.2 34.1$\pm$2.8 96.4$\pm$1.4 30.1$\pm$1.6 88.7$\pm$2.5 32.2 91.9
+8x 27.0$\pm$2.5 78.5$\pm$4.5 29.1$\pm$3.2 90.2$\pm$3.2 27.1$\pm$1.6 84.2$\pm$3.3 27.7 84.3
+FedGIMP 4x 33.2$\pm$2.3 93.1$\pm$1.5 31.7$\pm$3.9 94.3$\pm$4.3 30.1$\pm$2.1 88.6$\pm$2.4 31.7 92.0
+8x 29.2$\pm$2.1 83.6$\pm$3.1 28.8$\pm$3.1 90.7$\pm$5.2 26.3$\pm$2.7 82.7$\pm$3.5 28.1 85.7
+FedDDA 4x 32.5$\pm$2.3 91.8$\pm$1.8 32.7$\pm$3.5 94.7$\pm$3.4 24.5$\pm$2.6 81.2$\pm$3.5 29.9 89.2
+8x 28.1$\pm$2.2 81.9$\pm$2.9 29.6$\pm$3.3 90.9$\pm$5.5 18.4$\pm$2.0 68.7$\pm$5.7 25.4 80.5
+FedGAT 4x 34.1$\pm$2.2 97.3$\pm$1.7 35.9$\pm$2.8 96.9$\pm$1.2 31.9$\pm$1.7 90.2$\pm$2.5 34.0 94.8
+8x 29.5$\pm$2.2 83.7$\pm$3.1 31.5$\pm$2.3 92.9$\pm$2.4 28.3$\pm$1.9 85.2$\pm$3.6 29.8 87.3
+Across-site Site 1 (MoDL) Site 2 (rGAN) Site 3 (D5C5) Average
+[…]
+non-FL Single 4x 38.4$\pm$3.3 97.3$\pm$1.7 31.2$\pm$3.2 93.0$\pm$2.6 26.2$\pm$2.6 79.4$\pm$5.7 31.9 89.9
+8x 32.4$\pm$3.0 92.6$\pm$2.9 26.8$\pm$3.2 85.1$\pm$5.1 22.6$\pm$2.3 70.3$\pm$6.9 27.3 82.7
+Central 4x 39.7$\pm$3.4 97.6$\pm$1.7 37.4$\pm$2.1 96.3$\pm$1.3 27.5$\pm$2.5 81.6$\pm$5.4 34.9 91.8
+8x 34.3$\pm$3.1 93.9$\pm$3.0 34.0$\pm$2.1 91.9$\pm$2.8 25.1$\pm$2.5 74.4$\pm$6.5 31.1 86.7
+FL methods FedDF 4x 38.0$\pm$3.0 96.5$\pm$1.4 32.3$\pm$2.9 92.7$\pm$1.8 26.7$\pm$2.6 81.1$\pm$5.3 32.3 90.1
+8x 32.1$\pm$2.8 92.3$\pm$3.1 27.7$\pm$2.8 85.1$\pm$4.5 24.0$\pm$2.5 70.6$\pm$6.9 27.9 82.6
+FedMD 4x 37.8$\pm$3.1 96.5$\pm$1.6 32.6$\pm$2.9 92.7$\pm$1.7 26.4$\pm$2.5 81.0$\pm$5.3 32.3 90.1
+8x 31.4$\pm$2.9 91.2$\pm$3.2 28.0$\pm$2.5 84.4$\pm$4.7 24.0$\pm$2.4 70.1$\pm$6.6 27.8 81.9
+FedGIMP 4x 37.6$\pm$3.2 97.2$\pm$1.6 26.6$\pm$3.1 85.7$\pm$5.7 26.0$\pm$2.5 80.0$\pm$5.7 30.1 87.6
+8x 32.1$\pm$3.0 92.5$\pm$2.9 22.9$\pm$2.9 76.0$\pm$8.3 22.2$\pm$2.3 67.9$\pm$8.8 25.7 78.8
+FedDDA 4x 35.5$\pm$3.3 94.5$\pm$2.4 27.0$\pm$3.5 84.6$\pm$4.7 22.3$\pm$2.3 72.9$\pm$6.6 28.3 84.0
+8x 30.2$\pm$3.6 88.7$\pm$4.0 24.0$\pm$3.4 75.7$\pm$5.9 19.2$\pm$2.3 60.7$\pm$9.2 24.5 75.0
+FedGAT 4x 39.3$\pm$3.1 97.5$\pm$1.5 34.0$\pm$2.3 93.5$\pm$2.0 27.2$\pm$2.6 81.2$\pm$5.6 33.5 90.7
+8x 33.5$\pm$3.0 92.9$\pm$3.1 29.2$\pm$2.3 86.6$\pm$4.2 24.5$\pm$2.6 77.6$\pm$8.8 29.1 85.7
 4.3 Modeling Procedures
@@
 The reconstruction models trained in these setups included MoDL [13], rGAN [19], and D5C5 [12].
-All sites were taken to use common imaging operators with matched acceleration rates and k-space sampling densities.
+All sites used a common acceleration rate and k-space sampling density.
 Each site was allowed to train a distinct reconstruction model.
-In a first experiment, the three sites trained entirely distinct architectures (MoDL-3, rGAN, D5C5).
-In a second experiment, the sites trained the same model type albeit using different numbers of cascades (MoDL-3, MoDL-5, MoDL-7).
+In a first experiment, the three sites trained entirely distinct architectures based on adversarial or unrolled backbones (MoDL, rGAN, D5C5).
+In a second experiment, the sites trained the same type of physics-driven model, albeit using different numbers of cascades in the unrolled backbone (MoDL3, MoDL5, MoDL7).
+In experiments involving a non-participating site, an unrolled TransUNet [79] architecture was used.
 All models were implemented using the PyTorch framework and executed on RTX 4090 GPUs.
 Models were trained via the AdamW optimizer with $\beta_{1}$=\,=\,=0.9, $\beta_{2}$=\,=\,=0.95, a weight decay of 0.05, and a learning rate of $\eta$=\,=\,=$10^{-4}$.
-As model-heterogeneous settings preclude aggregation via simple averaging of model weights, FL baselines used either model distillation or synthetic image generation as a means for cross-site knowledge transfer.
+As model-heterogeneous settings preclude aggregation via simple weight averaging, FL baselines used either model distillation or synthetic image generation as a means for cross-site knowledge transfer.
 Note that distillation-based aggregation involves bidirectional distillation between global and site-specific models, with each direction incurring significantly higher computational costs compared to model averaging, as the student model must be trained based on a sizable set of input-output images produced by the teacher model.
@@
 Performance differences among competing methods were examined via non-parametric Wilcoxon signed-rank tests (p$<$0.05).
-Figure 3: Representative reconstructions at R=8x from zero-filled Fourier method (Zero-filled), single-site models (Single), FL baselines (FedDF, FedMD, FedGIMP, FedDDA), and FedGAT, along with reference images.
-(a) Site-specific BraTS model tested on fastMRI, (b) Site-specific fastMRI model tested on MIDAS, (c) Site-specific MIDAS model on BraTS.
-Zoom-in displays are included showing error maps (left) and images (right) within a focal window to emphasize method differences.
-Table 3: Within-site and cross-site reconstruction performance for single-coil datasets, where each site employs a variant of MoDL with different number of cascades.
-Within-site Site 1 (BraTS-MoDL-5) Site 2 (fastMRI-MoDL-3) Site 3 (MIDAS-MoDL-7)
-Reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
-Non-FL Single 4x 51.80$\pm$2.83 99.86$\pm$0.06 40.79$\pm$2.94 97.09$\pm$1.26 36.31$\pm$2.89 95.91$\pm$1.71
-8x 44.07$\pm$3.26 99.38$\pm$0.33 36.30$\pm$2.83 95.14$\pm$1.81 30.93$\pm$2.81 92.45$\pm$3.01
-Central 4x 51.30$\pm$2.38 99.83$\pm$0.08 40.00$\pm$3.12 96.76$\pm$1.36 36.06$\pm$2.76 95.73$\pm$1.73
-8x 44.19$\pm$2.79 99.26$\pm$0.36 35.02$\pm$3.00 94.14$\pm$2.05 31.08$\pm$2.80 92.47$\pm$2.97
-FL methods FedDF 4x 47.08$\pm$2.88 99.63$\pm$0.13 39.31$\pm$3.12 96.35$\pm$1.55 33.31$\pm$2.71 94.19$\pm$2.13
-8x 41.53$\pm$3.13 98.89$\pm$0.46 34.18$\pm$3.13 93.49$\pm$2.48 29.54$\pm$2.55 90.28$\pm$3.28
-FedMD 4x 47.02$\pm$2.84 99.63$\pm$0.13 39.23$\pm$3.27 96.37$\pm$1.58 33.37$\pm$2.69 94.20$\pm$2.12
-8x 41.43$\pm$3.10 98.86$\pm$0.47 33.97$\pm$3.16 93.52$\pm$2.46 30.03$\pm$2.53 90.52$\pm$3.25
-FedGIMP 4x 46.33$\pm$3.07 99.56$\pm$0.14 39.46$\pm$3.60 96.38$\pm$1.78 34.71$\pm$2.88 94.66$\pm$2.21
-8x 41.60$\pm$2.89 98.72$\pm$0.42 34.90$\pm$3.32 93.97$\pm$2.57 30.28$\pm$2.78 90.56$\pm$3.63
-FedDDA 4x 47.38$\pm$2.81 99.66$\pm$0.12 39.22$\pm$3.35 96.53$\pm$1.51 29.16$\pm$1.39 60.74$\pm$2.14
-8x 42.92$\pm$2.99 99.08$\pm$0.40 34.14$\pm$3.36 93.86$\pm$2.43 27.80$\pm$2.03 63.50$\pm$2.74
-FedGAT 4x 50.42$\pm$2.44 99.81$\pm$0.08 40.01$\pm$3.20 96.78$\pm$1.35 35.33$\pm$2.90 95.57$\pm$1.72
-8x 43.43$\pm$2.85 99.25$\pm$0.35 35.50$\pm$2.91 94.54$\pm$2.10 30.04$\pm$2.87 91.92$\pm$3.10
-Across-site Site 1 (BraTS-MoDL-5) Site 2 (fastMRI-MoDL-3) Site 3 (MIDAS-MoDL-7)
-Reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
-Non-FL Single 4x 34.21$\pm$4.48 90.92$\pm$4.36 39.25$\pm$6.94 95.83$\pm$4.14 45.38$\pm$5.04 98.28$\pm$1.77
-8x 30.79$\pm$4.08 87.83$\pm$5.52 34.46$\pm$6.48 92.76$\pm$5.57 38.99$\pm$4.18 97.09$\pm$2.44
-Central 4x 38.98$\pm$4.55 96.36$\pm$1.80 40.83$\pm$7.05 96.78$\pm$3.19 47.23$\pm$5.35 98.62$\pm$1.50
-8x 33.82$\pm$4.45 93.52$\pm$3.15 35.31$\pm$6.20 93.43$\pm$5.26 41.77$\pm$4.41 97.75$\pm$2.03
-FL methods FedDF 4x 36.36$\pm$2.96 95.21$\pm$1.92 39.33$\pm$2.66 96.23$\pm$1.74 43.44$\pm$3.03 98.09$\pm$1.04
-8x 31.94$\pm$2.87 92.09$\pm$2.86 34.91$\pm$2.80 93.62$\pm$2.48 37.86$\pm$3.15 96.31$\pm$1.75
-FedMD 4x 36.28$\pm$2.92 95.24$\pm$1.87 39.55$\pm$2.75 96.28$\pm$1.79 43.43$\pm$3.02 98.08$\pm$1.04
-8x 31.91$\pm$2.85 92.16$\pm$2.81 34.87$\pm$2.84 93.63$\pm$2.53 37.58$\pm$3.25 96.37$\pm$1.72
-FedGIMP 4x 36.97$\pm$4.56 94.87$\pm$3.09 38.71$\pm$6.51 95.25$\pm$4.62 44.10$\pm$4.08 98.33$\pm$1.74
-8x 32.89$\pm$4.50 92.33$\pm$3.89 34.65$\pm$6.20 92.31$\pm$5.76 39.02$\pm$3.74 96.84$\pm$2.79
-FedDDA 4x 37.52$\pm$4.02 95.48$\pm$2.51 39.81$\pm$6.80 95.91$\pm$4.06 32.27$\pm$1.40 59.96$\pm$6.01
-8x 32.98$\pm$4.21 92.71$\pm$3.53 35.06$\pm$6.15 93.16$\pm$5.36 31.45$\pm$1.78 59.85$\pm$7.96
-FedGAT 4x 38.07$\pm$4.48 96.09$\pm$1.93 40.32$\pm$7.04 96.33$\pm$3.66 46.68$\pm$5.21 98.68$\pm$1.65
-8x 33.35$\pm$4.23 93.16$\pm$3.37 35.13$\pm$6.94 93.64$\pm$5.43 40.65$\pm$4.37 97.42$\pm$2.31
-Table 4: Within-site and cross-site reconstruction performance for multi-coil datasets, where each site employs a variant of MoDL with different number of cascades.
-Within-site Site 1 (f.knee-MoDL-3) Site 2 (f.brain-MoDL-5) Site 3 (UMRAM-MoDL-7)
-Reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
-Non-FL Single 4x 33.78$\pm$2.27 93.37$\pm$1.52 40.33$\pm$3.45 98.03$\pm$0.84 49.96$\pm$2.06 99.61$\pm$0.17
-8x 29.39$\pm$2.06 83.90$\pm$3.08 34.21$\pm$3.36 94.33$\pm$2.01 42.11$\pm$1.90 98.25$\pm$0.73
-Central 4x 33.69$\pm$2.32 93.20$\pm$1.63 40.95$\pm$3.84 98.05$\pm$1.19 46.36$\pm$2.22 99.32$\pm$0.30
-8x 29.39$\pm$2.19 83.62$\pm$3.09 34.27$\pm$3.50 94.11$\pm$2.28 39.73$\pm$1.86 97.87$\pm$0.72
-FL methods FedDF 4x 32.33$\pm$2.36 90.47$\pm$2.44 38.87$\pm$3.62 95.66$\pm$1.49 44.56$\pm$2.72 98.52$\pm$0.63
-8x 28.36$\pm$2.31 80.66$\pm$4.14 32.97$\pm$3.73 91.37$\pm$3.37 37.00$\pm$2.30 96.73$\pm$1.40
-FedMD 4x 32.93$\pm$2.35 91.96$\pm$1.77 39.94$\pm$3.94 97.49$\pm$1.09 44.63$\pm$3.05 99.25$\pm$0.28
-8x 28.15$\pm$2.42 81.60$\pm$3.52 33.85$\pm$3.79 93.68$\pm$2.55 37.55$\pm$2.55 97.09$\pm$1.20
-FedGIMP 4x 33.78$\pm$2.27 93.41$\pm$1.48 40.54$\pm$3.38 97.88$\pm$0.90 44.99$\pm$2.09 99.02$\pm$0.32
-8x 28.97$\pm$2.12 83.68$\pm$3.07 34.19$\pm$3.37 93.80$\pm$3.22 39.53$\pm$1.70 97.68$\pm$0.77
-FedDDA 4x 32.53$\pm$2.33 91.75$\pm$1.82 39.60$\pm$3.62 97.52$\pm$1.38 42.57$\pm$2.92 97.32$\pm$0.78
-8x 28.09$\pm$2.22 81.90$\pm$2.93 33.25$\pm$3.52 93.61$\pm$2.76 37.51$\pm$2.17 93.63$\pm$1.50
-FedGAT 4x 34.10$\pm$2.24 93.44$\pm$1.46 40.87$\pm$3.58 98.03$\pm$1.27 45.13$\pm$2.14 99.22$\pm$0.29
-8x 29.49$\pm$2.21 83.73$\pm$3.07 34.20$\pm$3.48 93.85$\pm$2.52 40.21$\pm$1.71 97.84$\pm$0.66
-Across-site Site 1 (f.knee-MoDL-3) Site 2 (f.brain-MoDL-5) Site 3 (UMRAM-MoDL-7)
-Reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
-Non-FL Single 4x 38.40$\pm$3.26 97.32$\pm$1.67 39.42$\pm$4.49 96.91$\pm$2.13 37.60$\pm$4.58 96.03$\pm$2.34
-8x 32.39$\pm$3.04 92.63$\pm$2.94 32.99$\pm$3.57 90.55$\pm$6.10 31.48$\pm$3.96 88.07$\pm$5.35
-Central 4x 39.70$\pm$3.36 97.62$\pm$1.70 40.64$\pm$4.84 97.36$\pm$1.99 40.39$\pm$4.11 97.53$\pm$1.31
-8x 34.33$\pm$3.12 93.86$\pm$3.04 33.61$\pm$3.89 91.38$\pm$5.74 34.47$\pm$3.50 91.89$\pm$4.04
-FL methods FedDF 4x 38.15$\pm$3.41 96.40$\pm$2.11 38.28$\pm$4.64 95.18$\pm$3.56 37.37$\pm$4.12 94.73$\pm$2.39
-8x 32.94$\pm$3.12 92.43$\pm$3.47 32.69$\pm$4.11 89.39$\pm$7.34 30.15$\pm$3.64 86.84$\pm$5.61
-FedMD 4x 38.75$\pm$3.44 97.01$\pm$1.71 39.32$\pm$4.78 96.80$\pm$2.28 38.39$\pm$4.49 95.92$\pm$2.28
-8x 32.52$\pm$2.90 92.24$\pm$3.02 33.19$\pm$3.76 90.90$\pm$6.08 31.90$\pm$3.81 87.51$\pm$5.55
-FedGIMP 4x 37.87$\pm$3.13 97.23$\pm$1.50 38.93$\pm$4.32 96.75$\pm$2.16 37.40$\pm$4.38 96.35$\pm$1.76
-8x 32.01$\pm$2.98 92.52$\pm$2.91 32.96$\pm$3.63 90.71$\pm$5.94 31.70$\pm$4.13 89.29$\pm$5.18
-FedDDA 4x 35.54$\pm$3.34 94.49$\pm$2.44 37.78$\pm$3.78 95.65$\pm$1.52 36.61$\pm$4.16 95.15$\pm$2.36
-8x 30.15$\pm$3.57 88.74$\pm$3.98 32.10$\pm$3.43 89.21$\pm$5.03 30.80$\pm$3.74 86.32$\pm$5.53
-FedGAT 4x 39.25$\pm$3.09 97.50$\pm$1.52 39.65$\pm$4.53 97.16$\pm$2.07 39.16$\pm$4.33 97.10$\pm$1.60
-8x 33.52$\pm$2.98 92.90$\pm$3.11 33.58$\pm$3.93 90.93$\pm$5.81 32.98$\pm$3.89 90.29$\pm$4.57
 4.4 Datasets
-In single-coil experiments, multi-contrast coil-combined magnitude MR images from public BraTS [70], fastMRI [71] and MIDAS [72] datasets were analyzed.
+In coil-combined experiments, multi-contrast coil-combined magnitude MR images from public BraTS [80], fastMRI [81] and MIDAS [82] datasets were analyzed.
 Each dataset was taken to correspond to a separate site.
@@
 Experimental procedures for collecting the in-house dataset were approved by the local ethics committee at Bilkent University, and written informed consent was obtained from all participants.
+In experiments involving a non-participating site, multi-coil k-space data from the Calgary dataset [77] were analyzed.
 Each dataset was taken to correspond to a separate site.
@@
 Multi-coil data were compressed to 5 virtual coils [83].
-Both actual and synthetic k-space data were retrospectively undersampled using a variable-density (VD) pattern with acceleration factors R = 4x and 8x [1].
-Models were trained on a mixture of multi-contrast MRI datasets without special procedures to handle separate contrasts.
-When necessary, actual and synthetic MR images were zero-padded to a consistent size to permit training of reconstruction models, albeit performance assessments were conducted in native image sizes by discarding the padded regions following reconstruction.
+Both actual and synthetic k-space data were retrospectively undersampled using a variable-density (VD) pattern with acceleration factors R = 4x-8x [1].
+Models were trained on a mixture of multi-contrast MRI datasets.
+When necessary, actual and synthetic MR images were zero-padded to a consistent size to permit training of reconstruction models, and performance assessments were conducted in native image sizes by discarding the padded regions following reconstruction.
+Figure 3: Representative reconstructions at R=8x from zero-filled Fourier method (Zero-filled), single-site models (Single), FL baselines (FedDF, FedMD, FedGIMP, FedDDA), and FedGAT, along with reference images.
+(a) Site-specific BraTS model tested on fastMRI, (b) Site-specific fastMRI model tested on MIDAS, (c) Site-specific MIDAS model tested on BraTS.
+Zoom-in displays are included showing error maps (left; see colorbar) and images (right) within a focal window to emphasize method differences.
+Table 3: Within-site and across-site reconstruction performance for coil-combined datasets (Site 1: BraTS, Site 2: fastMRI, Site 3: MIDAS).
+Each site employs an MoDL variant with different number of cascades.
+Within-site Site 1 (MoDL5) Site 2 (MoDL3) Site 3 (MoDL7) Average
+[…]
+Non-FL Single 4x 51.8$\pm$2.8 99.9$\pm$0.1 40.8$\pm$2.9 97.1$\pm$1.3 36.3$\pm$2.9 95.9$\pm$1.7 43.0 97.6
+8x 44.1$\pm$3.3 99.4$\pm$0.3 36.3$\pm$2.8 95.1$\pm$1.8 30.9$\pm$2.8 92.5$\pm$3.0 37.1 95.7
+Central 4x 51.3$\pm$2.4 99.8$\pm$0.1 40.0$\pm$3.1 96.8$\pm$1.4 36.1$\pm$2.8 95.7$\pm$1.7 42.5 97.4
+8x 44.2$\pm$2.8 99.3$\pm$0.4 35.0$\pm$3.0 94.1$\pm$2.0 31.1$\pm$2.8 92.5$\pm$3.0 36.8 95.3
+FL methods FedDF 4x 47.1$\pm$2.9 99.6$\pm$0.1 39.3$\pm$3.1 96.4$\pm$1.6 33.3$\pm$2.7 94.2$\pm$2.1 39.9 96.7
+8x 41.5$\pm$3.1 98.9$\pm$0.5 34.2$\pm$3.1 93.5$\pm$2.5 29.5$\pm$2.6 90.3$\pm$3.3 35.1 94.2
+FedMD 4x 47.0$\pm$2.8 99.6$\pm$0.1 39.2$\pm$3.3 96.4$\pm$1.6 33.4$\pm$2.7 94.2$\pm$2.1 39.9 96.7
+8x 41.4$\pm$3.1 98.9$\pm$0.5 34.0$\pm$3.2 93.5$\pm$2.5 30.0$\pm$2.5 90.5$\pm$3.3 35.1 94.3
+FedGIMP 4x 46.3$\pm$3.1 99.6$\pm$0.1 39.5$\pm$3.6 96.4$\pm$1.8 34.7$\pm$2.9 94.7$\pm$2.2 40.2 96.9
+8x 41.6$\pm$2.9 98.7$\pm$0.4 34.9$\pm$3.3 94.0$\pm$2.6 30.3$\pm$2.8 90.6$\pm$3.6 35.6 94.4
+FedDDA 4x 47.4$\pm$2.8 99.7$\pm$0.1 39.2$\pm$3.4 96.5$\pm$1.5 29.2$\pm$1.4 60.7$\pm$2.1 38.6 85.6
+8x 42.9$\pm$3.0 99.1$\pm$0.4 34.1$\pm$3.4 93.9$\pm$2.4 27.8$\pm$2.0 63.5$\pm$2.7 34.9 85.5
+FedGAT 4x 50.4$\pm$2.4 99.8$\pm$0.1 39.8$\pm$3.2 96.7$\pm$1.3 36.0$\pm$2.9 95.8$\pm$1.8 42.1 97.4
+8x 43.4$\pm$2.9 99.3$\pm$0.4 35.3$\pm$3.1 94.6$\pm$2.1 30.6$\pm$2.8 92.0$\pm$3.1 36.4 95.3
+Across-site Site 1 (MoDL5) Site 2 (MoDL3) Site 3 (MoDL7) Average
+[…]
+Non-FL Single 4x 34.2$\pm$4.5 90.9$\pm$4.4 39.3$\pm$6.9 95.8$\pm$4.1 45.4$\pm$5.0 98.3$\pm$1.8 39.6 95.0
+8x 30.8$\pm$4.1 87.8$\pm$5.5 34.5$\pm$6.5 92.8$\pm$5.6 39.0$\pm$4.2 97.1$\pm$2.4 34.8 92.6
+Central 4x 39.0$\pm$4.6 96.4$\pm$1.8 40.8$\pm$7.1 96.8$\pm$3.2 47.2$\pm$5.4 98.6$\pm$1.5 42.3 97.3
+8x 33.8$\pm$4.5 93.5$\pm$3.2 35.3$\pm$6.2 93.4$\pm$5.3 41.8$\pm$4.4 97.8$\pm$2.0 37.0 94.9
+FL methods FedDF 4x 36.4$\pm$3.0 95.2$\pm$1.9 39.3$\pm$2.7 96.2$\pm$1.7 43.4$\pm$3.0 98.1$\pm$1.0 39.7 96.5
+8x 31.9$\pm$2.9 92.1$\pm$2.9 34.9$\pm$2.8 93.6$\pm$2.5 37.9$\pm$3.2 96.3$\pm$1.8 34.9 94.0
+FedMD 4x 36.3$\pm$2.9 95.2$\pm$1.9 39.6$\pm$2.8 96.3$\pm$1.8 43.4$\pm$3.0 98.1$\pm$1.0 39.8 96.5
+8x 31.9$\pm$2.9 92.2$\pm$2.8 34.9$\pm$2.8 93.6$\pm$2.5 37.6$\pm$3.3 96.4$\pm$1.7 34.8 94.1
+FedGIMP 4x 37.0$\pm$4.6 94.9$\pm$3.1 38.7$\pm$6.5 95.3$\pm$4.6 44.1$\pm$4.1 98.3$\pm$1.7 39.9 96.2
+8x 32.9$\pm$4.5 92.3$\pm$3.9 34.7$\pm$6.2 92.3$\pm$5.8 39.0$\pm$3.7 96.8$\pm$2.8 35.5 93.8
+FedDDA 4x 37.5$\pm$4.0 95.5$\pm$2.5 39.8$\pm$6.8 95.9$\pm$4.1 32.3$\pm$1.4 60.0$\pm$6.0 36.5 83.8
+8x 33.0$\pm$4.2 92.7$\pm$3.5 35.1$\pm$6.2 93.2$\pm$5.4 31.5$\pm$1.8 59.9$\pm$8.0 33.2 81.9
+FedGAT 4x 38.1$\pm$4.5 96.1$\pm$1.9 40.1$\pm$7.0 96.4$\pm$3.6 46.8$\pm$5.3 98.4$\pm$1.9 41.7 97.0
+8x 33.4$\pm$4.2 93.2$\pm$3.4 35.3$\pm$6.8 93.8$\pm$5.4 40.4$\pm$4.6 97.4$\pm$2.4 36.4 94.8
+Table 4: Within-site and across-site reconstruction performance for multi-coil datasets (Site 1: fastMRI-knee, Site 2: fastMRI-brain, Site 3: UMRAM).
+Each site employs an MoDL variant with different number of cascades.
+Within-site Site 1 (MoDL3) Site 2 (MoDL5) Site 3 (MoDL7) Average
+[…]
+Non-FL Single 4x 33.8$\pm$2.3 93.4$\pm$1.5 40.3$\pm$3.5 98.0$\pm$0.8 50.0$\pm$2.1 99.6$\pm$0.2 41.4 97.0
+8x 29.4$\pm$2.1 83.9$\pm$3.1 34.2$\pm$3.4 94.3$\pm$2.0 42.1$\pm$1.9 98.3$\pm$0.7 35.2 92.2
+Central 4x 33.7$\pm$2.3 93.2$\pm$1.6 41.0$\pm$3.8 98.1$\pm$1.2 46.4$\pm$2.2 99.3$\pm$0.3 40.4 96.9
+8x 29.4$\pm$2.2 83.6$\pm$3.1 34.3$\pm$3.5 94.1$\pm$2.3 39.7$\pm$1.9 97.9$\pm$0.7 34.5 91.9
+FL methods FedDF 4x 32.3$\pm$2.4 90.5$\pm$2.4 38.9$\pm$3.6 95.7$\pm$1.5 44.6$\pm$2.7 98.5$\pm$0.6 38.6 94.9
+8x 28.4$\pm$2.3 80.7$\pm$4.1 33.0$\pm$3.7 91.4$\pm$3.4 37.0$\pm$2.3 96.7$\pm$1.4 32.8 89.6
+FedMD 4x 32.9$\pm$2.4 92.0$\pm$1.8 39.9$\pm$3.9 97.5$\pm$1.1 44.6$\pm$3.1 99.2$\pm$0.3 39.1 96.3
+8x 28.2$\pm$2.4 81.6$\pm$3.5 33.9$\pm$3.8 93.7$\pm$2.6 37.6$\pm$2.6 97.1$\pm$1.2 33.2 90.8
+FedGIMP 4x 33.2$\pm$2.3 93.1$\pm$1.5 40.5$\pm$3.4 97.9$\pm$0.9 45.0$\pm$2.1 99.0$\pm$0.3 39.6 96.7
+8x 29.2$\pm$2.1 83.6$\pm$3.1 33.5$\pm$3.6 93.8$\pm$2.3 39.5$\pm$1.7 97.7$\pm$0.8 34.1 91.7
+FedDDA 4x 32.5$\pm$2.3 91.8$\pm$1.8 39.6$\pm$3.6 97.5$\pm$1.4 42.6$\pm$2.9 97.3$\pm$0.8 38.2 95.5
+8x 28.1$\pm$2.2 81.9$\pm$2.9 33.3$\pm$3.5 93.6$\pm$2.8 37.5$\pm$2.2 93.6$\pm$1.5 33.0 89.7
+FedGAT 4x 34.1$\pm$2.2 97.3$\pm$1.7 40.9$\pm$3.6 98.0$\pm$1.3 45.1$\pm$2.1 99.2$\pm$0.3 40.0 96.9
+8x 29.5$\pm$2.2 83.7$\pm$3.1 34.2$\pm$3.5 93.9$\pm$2.5 40.2$\pm$1.7 97.8$\pm$0.7 34.6 91.8
+Across-site Site 1 (MoDL3) Site 2 (MoDL5) Site 3 (MoDL7) Average
+[…]
+Non-FL Single 4x 38.4$\pm$3.3 97.3$\pm$1.7 39.4$\pm$4.5 96.9$\pm$2.1 37.6$\pm$4.6 96.0$\pm$2.3 38.5 96.7
+8x 32.4$\pm$3.0 92.6$\pm$2.9 33.0$\pm$3.6 90.6$\pm$6.1 31.5$\pm$4.0 88.1$\pm$5.4 32.3 90.4
+Central 4x 39.7$\pm$3.4 97.6$\pm$1.7 40.6$\pm$4.8 97.4$\pm$2.0 40.4$\pm$4.1 97.5$\pm$1.3 40.2 97.5
+8x 34.3$\pm$3.1 93.9$\pm$3.0 33.6$\pm$3.9 91.4$\pm$5.7 34.5$\pm$3.5 91.9$\pm$4.0 34.1 92.4
+FL methods FedDF 4x 38.2$\pm$3.4 96.4$\pm$2.1 38.3$\pm$4.6 95.2$\pm$3.6 37.4$\pm$4.1 94.7$\pm$2.4 38.0 95.4
+8x 32.9$\pm$3.1 92.4$\pm$3.5 32.7$\pm$4.1 89.4$\pm$7.3 30.2$\pm$3.6 86.8$\pm$5.6 31.9 89.5
+FedMD 4x 38.8$\pm$3.4 97.0$\pm$1.7 39.3$\pm$4.8 96.8$\pm$2.3 38.4$\pm$4.5 95.9$\pm$2.3 38.8 96.6
+8x 32.5$\pm$2.9 92.2$\pm$3.0 33.2$\pm$3.8 90.9$\pm$6.1 31.9$\pm$3.8 87.5$\pm$5.6 32.5 90.2
+FedGIMP 4x 37.6$\pm$3.2 97.2$\pm$1.6 38.9$\pm$4.3 96.8$\pm$2.2 37.4$\pm$4.4 96.4$\pm$1.8 38.0 96.8
+8x 32.1$\pm$3.0 92.5$\pm$2.9 32.5$\pm$3.5 90.6$\pm$6.1 31.7$\pm$4.1 89.3$\pm$5.2 32.1 90.8
+FedDDA 4x 35.5$\pm$3.3 94.5$\pm$2.4 37.8$\pm$3.8 95.7$\pm$1.5 36.6$\pm$4.2 95.2$\pm$2.4 36.6 95.1
+8x 30.2$\pm$3.6 88.7$\pm$4.0 32.1$\pm$3.4 89.2$\pm$5.0 30.8$\pm$3.7 86.3$\pm$5.5 31.0 88.1
+FedGAT 4x 39.3$\pm$3.1 97.5$\pm$1.5 39.7$\pm$4.5 97.2$\pm$2.1 39.2$\pm$4.3 97.1$\pm$1.6 39.4 97.3
+8x 33.5$\pm$3.0 92.9$\pm$3.1 33.6$\pm$3.9 90.9$\pm$5.8 33.0$\pm$3.9 90.3$\pm$4.6 33.4 91.4
 5 Results
@@
 ‘Central’ models trained on the server using multi-site data aggregated across all sites were taken as a privacy-violating performance benchmark for across-site reconstructions.
-In a first FL setup, we examined performance when individual sites employed entirely distinct types of model architectures (i.e., MoDL-3, rGAN, D5C5).
-Performance metrics are listed in Table 1 for single-coil and in Table 2 for multi-coil MRI datasets.
-For within-site reconstructions, FedGAT achieves significantly higher performance than all competing FL methods (p$<$0.05), except for FedGIMP that yields higher PSNR for R=8x on single-coil data at Site 1.
-On average, FedGAT offers (PSNR, SSIM) improvements of (2.06dB, 4.40%) over distillation baselines, and (2.07dB, 2.48)SSIM over generative baselines.
-Among FL methods, FedGAT attains the most competitive performance against the ‘Single’ benchmark, and outperforms the privacy-violating ‘Central’ by 0.42dB PSNR, 0.71% SSIM (p$<$0.05).
+In a first FL setup, we examined performance when individual sites employed entirely distinct types of model architectures (i.e., MoDL, rGAN, D5C5).
+Performance metrics are listed in Table 1 for coil-combined and in Table 2 for multi-coil MRI datasets.
+For within-site reconstructions, FedGAT achieves significantly higher performance than all competing FL methods (p$<$0.05), except for FedGIMP that yields higher PSNR for R=8x on coil-combined data at Site 1.
+On average, FedGAT offers (PSNR, SSIM) improvements of (2.1dB, 4.4%) over distillation baselines, and (2.1dB, 2.5%) over generative baselines.
+Among FL methods, FedGAT attains the most competitive performance against the ‘Single’ benchmark, and outperforms the privacy-violating ‘Central’ by 0.5dB PSNR, 0.7% SSIM (p$<$0.05).
 For across-site reconstructions, FedGAT achieves significantly higher performance than all competing FL methods (p$<$0.05).
-On average, FedGAT offers (PSNR, SSIM) improvements of (1.22dB, 3.48%) over distillation baselines, and (2.60dB, 5.74)SSIM over generative baselines.
-In this setting, FedGAT attains the most competitive performance against the ‘Central’ benchmark while outperforming ‘Single’ by 1.59dB PSNR, 4.20% SSIM (p$<$0.05), which showcasing the generalization benefits of training models on multi-site datasets.
+On average, FedGAT offers (PSNR, SSIM) improvements of (1.2dB, 3.5%) over distillation baselines, and (2.6dB, 5.7%) over generative baselines.
+In this setting, FedGAT attains the most competitive performance against the ‘Central’ benchmark while outperforming ‘Single’ by 1.6dB PSNR, 4.2% SSIM (p$<$0.05), which showcases the generalization benefits of training models on multi-site datasets.
 Representative images reconstructed via FL methods in the examined model-heterogeneous setup are depicted in Fig.
 3.
-Corroborating the quantitative assessments, FedGAT attains higher image quality compared to baselines, with sharper depiction of structural details, lower artifact and noise in reconstructed images.
-Meanwhile, competing FL methods suffer from a degree of spatial blur, residual noise and artifacts that lead to suboptimal depiction of detailed tissue structures.
-In a second FL setup, we examined performance when individual sites employed distinct variants of a common model type (i.e., MoDL with 3, 5, versus 7 cascades), resulting in varying levels of model complexity across sites.
-Naturally, the benefits of model-agnostic approaches can be lower in this setup, when compared against the first setup that prescribed entirely distinct model types.
+Corroborating the quantitative assessments, FedGAT attains higher image quality compared to baselines, with sharper depiction of structural details, lower artifact and noise levels in reconstructed images.
+Meanwhile, competing FL methods suffer from a degree of spatial blur, residual noise or artifacts that lead to suboptimal depiction of detailed tissue structures.
+In a second FL setup, we examined performance when individual sites employed distinct variants of a common model type (i.e., the unrolled MoDL model with 3, 5, versus 7 cascades), reflecting a scenario in which sites prefer varying levels of model complexity.
+Naturally, the benefits of model-agnostic approaches can be moderately lower in this setup, when compared against the first setup that prescribed entirely distinct model types.
 Still, we reasoned that FedGAT should continue to yield significant benefits over competing FL baselines.
-Performance metrics are listed in Table 3 for single-coil and in Table 4 for multi-coil MRI datasets.
+Performance metrics are listed in Table 3 for coil-combined and in Table 4 for multi-coil MRI datasets.
 For within-site reconstructions, FedGAT achieves significantly higher performance than all competing FL methods, and attains the most competitive performances against the ‘Single’ and privacy-violating ‘Central’ (p$<$0.05), except for FedMD that yields higher PSNR on multi-coil data at Site 3.
-On average, FedGAT offers (PSNR, SSIM) improvements of (1.52dB, 1.15%) over distillation baselines, and (1.42dB, 3.32)SSIM over generative baselines.
+On average, FedGAT offers (PSNR, SSIM) improvements of (1.6dB, 1.2%) over distillation baselines, and (1.5dB, 3.4%) over generative baselines.
 For across-site reconstructions, FedGAT achieves significantly higher performance than all competing FL methods (p$<$0.05).
-On average, FedGAT offers (PSNR, SSIM) improvements of (1.40dB, 0.99%) over distillation baselines, and (2.31dB, 4.29)SSIM over generative baselines.
-In this setting, FedGAT attains the most competitive performance against ‘Central’ while outperforming ‘Single’ by 1.42dB PSNR, 1.42% SSIM (p$<$0.05), indicating improved generalization.
+On average, FedGAT offers (PSNR, SSIM) improvements of (1.4dB, 1.0%) over distillation baselines, and (2.3dB, 4.3%) over generative baselines.
+In this setting, FedGAT attains the most competitive performance against ‘Central’ while outperforming ‘Single’ by 1.4dB PSNR, 1.4% SSIM (p$<$0.05), indicating improved generalization.
 Representative images reconstructed in the second model-heterogeneous FL setup are depicted in Fig.
@@
 Zoom-in windows of error maps and images are included to emphasize method differences.
-5.2 Ablation Studies
-The primary aim of the current study is to enable collaborative training of generalizable reconstruction models with heterogeneous architectures across imaging sites.
-Towards this aim, FedGAT leverages synthetic MR images produced by a generative autoregressive transformer (GAT) as a privacy-preserving means of cross-site knowledge transfer.
-To examine the importance of the proposed GAT prior, we compared the fidelity of synthetic MR images generated by the GAT prior against those generated by adversarial (GAN) or diffusion (DDPM) priors.
-Fidelity of multi-site MR images generated by different priors as measured by FID are displayed in Fig.
-5, and sample synthetic MR images are shown in Fig.
-6.
-We find that GAT achieves the highest image fidelity, as manifested by its consistently lower FID scores compared to variants, and by its improved visual realism apparent in sample images.
-While GAN shows structural artifacts and suboptimal contrast and DDPM occasionally shows elevated noise levels, GAT achieves high image quality with lower artifacts and noise.
-We then assessed whether this improvement in synthetic image fidelity translates onto benefits in reconstruction performance.
-For this purpose, performance of reconstruction models trained on synthetic images from the GAT, GAN and DDPM priors were compared as listed in Table 5, in reference to a benchmark model that was trained using actual multi-site MR images.
-We find that the GAT prior yields significantly higher reconstruction performance than variants based on GAN and DDPM priors (p$<$0.05), and it attains the most competitive performance against the benchmark model.
-These results corroborate that the superior image fidelity of the proposed GAT prior translates onto improving the training of site-specific reconstruction models.
-Next, we examined the efficacy of the knowledge-transfer approach in FedGAT that transfers information on the distribution of multi-site data via synthetic MR images.
-For this purpose, FedGAT was compared against a ‘w model agg.’ variant that simultaneously transferred knowledge and built reconstruction models via conventional model aggregation [36].
-The comparisons were performed in model-homogeneous settings to permit use of model aggregation, and to avoid interactions between model architecture and data distribution across sites that could bias assessments.
-Performance metrics are listed in Table 6.
-FedGAT yields superior performance against ‘w model agg.’ in all cases ((p$<$0.05), except for within-site reconstructions at Site 3 where ‘w model agg.’ yields slightly higher SSIM.
-These results indicate that the decoupling of knowledge transfer and model building stages in FedGAT offers performance benefits against a coupled model aggregation procedure.
-6 Discussion
-Here we introduced a model-agnostic FL technique for MRI reconstruction that supports model-heterogeneous settings by decoupling cross-site knowledge transfer from building of reconstruction models.
-In particular, FedGAT mediates knowledge transfer among sites by sharing a global GAT prior that uses autoregressive predictions across multiple spatial scales to generate synthetic MR images, and a site prompt for precise control over the synthetic image distribution.
-Previous FL studies aiming to build generative priors have predominantly adopted adversarial approaches that can suffer from poor image quality due to training instabilities [20, 18, 74], or diffusion approaches that can suffer from extensive inference times and residual image noise [75, 76, 77, 78].
-In comparison, the autoregressive nature of GAT offers improved image fidelity over adversarial and diffusion priors as indicated by our results.
-To curate a hybrid multi-site MRI dataset, the shared GAT prior is employed at each site to efficiently generate high-fidelity synthetic MRI datasets from remaining sites.
-Afterwards, site-specific reconstruction models are locally trained on this hybrid dataset to improve generalization while enjoying total freedom in architecture selection.
-Therefore, the proposed method holds great promise for advancing the practicality and scalability of FL in multi-institutional MRI studies.
+Figure 5: Performance of site-specific reconstruction models (TransUNet) trained at a held-out site (Calgary) using a mixture of local and synthetic datasets.
+Synthetic datasets were generated by a GAT prior built in a three-site FL setup (BraTS, fastMRI, MIDAS).
+Separate models were trained while the percentage of samples from the local dataset was systematically varied.
+Results plotted for within-site reconstructions at the held-out site (Calgary), as well as across-site reconstructions in the FL sites (see legend).
 Figure 6: Fidelity of synthetic MR images generated via GAN, DDPM and GAT priors. Bar plots of Frechet inception distance (FID) show mean±$\pm$±std across the test set.
 Figure 7: Samples of synthetic MR images from GAN, DDPM and GAT priors, along with samples of local MR images at separate sites.
+5.2 Generalization to a Non-Participating Site
+Our comparison studies indicate that FedGAT enables the development of site-specific reconstruction models that generalize reliably across the participating sites in model-heterogeneous FL setups.
+Yet, FedGAT can also be deployed at a non-participating site that builds its reconstruction model by benefiting from the aggregated knowledge encapsulated by the GAT prior, enhancing within-site performance in scenarios where only limited local training data are available, and promoting generalization to datasets that diverge from the local data distribution.
+To assess potential benefits of FedGAT in a non-participating site, we considered a three-site FL setup to build the GAT prior (BraTS, fastMRI, MIDAS), and deployed it to train a transformer-based reconstruction model (TransUNet) at a held-out site (Calgary).
+To systematically examine the influence of local dataset size at the held-out site, separate models were trained using varying proportions of local samples and synthetic samples generated by the GAT prior.
+For reference, a 0% local / 100% synthetic setting (denoted 0%) indicates training solely on synthetic data, while a 100% local / 0% synthetic setting (denoted 100%) indicates training solely on local data; intermediate percentages reflect mixtures of the two sources.
+Models trained at the held-out site were used to reconstruct images at the held-out site and also at each individual participating site, as shown in Fig.
+5.
+For within-site reconstructions at the held-out site, while the model trained only on synthetic samples (0%) moderately underperforms compared to the reference model trained only on local samples (100%), models trained with mixed data perform competitively.
+On average, the performance gap between the 0% model and the reference 100% model is -0.8dB in PSNR and -6.2% in SSIM.
+In contrast, models trained with 40%–80% local data exhibit average differences of only +0.2dB in PSNR and +0.7% in SSIM relative to the 100% model.
+This result indicates that synthetic MR images produced by FedGAT can support training of performant reconstruction models at non-participating sites even when local training data are scarce.
+For across-site reconstructions at participating sites, the model trained only on local samples (100%) exhibits poor generalization performance relative to the model trained only on synthetic samples (0%).
+However, models trained with a mixture of local and synthetic data (i.e., 20%–80% local) show improved generalization, with an average gain of 6.5dB in PSNR and 19.6% in SSIM compared to the 100% model.
+This finding suggests that incorporating synthetic samples can help boost the generalization capability of models trained at non-participating sites to other imaging sites.
 Table 5: Reconstruction performance of models trained on synthetic MR images generated by GAN, DDPM, and GAT priors, in reference to a benchmark model trained on actual MR images.
@@
 Within-site Site 1 (BraTS) Site 2 (fastMRI) Site 3 (MIDAS)
-reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
+[…]
 Actual 4x 47.38$\pm$2.61 99.61$\pm$0.15 40.00$\pm$3.12 96.76$\pm$1.36 34.29$\pm$2.63 93.95$\pm$2.06
@@
 Across-site Site 1 (BraTS) Site 2 (fastMRI) Site 3 (MIDAS)
-reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
+[…]
 Actual 4x 37.14$\pm$4.06 95.35$\pm$2.24 40.83$\pm$7.05 96.78$\pm$3.19 43.69$\pm$4.68 98.18$\pm$1.73
@@
 8x 32.11$\pm$4.09 91.53$\pm$3.95 35.22$\pm$6.67 93.72$\pm$5.46 38.20$\pm$4.29 96.40$\pm$2.72
-Table 6: Ablation study on the efficacy of the knowledge transfer approach in FedGAT.
+Table 6: Susceptibility to information leakage for GAN, DDPM and GAT priors.
+FID scores of synthetic images were measured in the absence of (w/o int) and in the presence of (w int) an intruder site in the FL setup.
+$\Delta$(w/o int - w int) denotes the degradation level in image quality due to the intruder site.
+Site 1 (BraTS) Site 2 (fastMRI) Site 3 (MIDAS) Average
+w/o int w int w/o int w int w/o int w int $\Delta$(w/o int - w int)
+GAN 7.47 14.05 1.15 2.77 2.68 6.43 3.98
+DDPM 10.55 16.02 6.02 11.36 8.92 11.87 4.59
+GAT 0.09 0.12 0.34 1.13 0.39 0.66 0.36
+Table 7: Reconstruction performance of FedGAT variants built to examine the contribution of individual method components. Boldface indicates the top-performing variant in each task.
 Within-site Site 1 (BraTS) Site 2 (fastMRI) Site 3 (MIDAS)
-reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
-FedGAT 4x 47.63$\pm$2.65 99.64$\pm$0.14 39.73$\pm$3.31 96.46$\pm$1.56 33.59$\pm$2.63 93.53$\pm$2.10
-8x 41.60$\pm$2.85 98.78$\pm$0.54 34.58$\pm$3.36 94.14$\pm$2.33 28.85$\pm$2.62 87.75$\pm$3.34
-w model agg. 4x 46.65$\pm$2.86 99.58$\pm$0.14 38.52$\pm$3.62 95.50$\pm$2.32 33.14$\pm$2.52 92.80$\pm$2.75
+[…]
+FedGAT 4x 47.90$\pm$2.70 99.66$\pm$0.13 39.76$\pm$3.15 96.73$\pm$1.34 34.02$\pm$2.62 93.72$\pm$2.18
+8x 41.61$\pm$2.88 98.66$\pm$0.60 35.34$\pm$3.05 94.60$\pm$2.12 29.56$\pm$2.53 89.35$\pm$3.29
+w one-hot 4x 44.48$\pm$2.64 98.86$\pm$0.56 39.24$\pm$2.47 96.33$\pm$1.40 33.21$\pm$2.39 92.48$\pm$2.13
+prompt 8x 38.00$\pm$2.64 87.56$\pm$5.25 34.35$\pm$2.58 92.30$\pm$1.87 28.96$\pm$2.47 85.56$\pm$3.07
+w/o site 4x 46.70$\pm$2.95 99.53$\pm$0.18 38.99$\pm$3.79 95.95$\pm$2.03 33.45$\pm$2.67 92.83$\pm$2.73
+prompt 8x 41.11$\pm$2.99 98.22$\pm$0.91 34.10$\pm$3.71 92.93$\pm$3.23 29.39$\pm$2.55 88.48$\pm$3.72
+w/o scales 4x 45.66$\pm$2.89 99.43$\pm$0.20 39.12$\pm$3.20 96.49$\pm$1.42 33.30$\pm$2.66 93.13$\pm$2.11
+8x 41.17$\pm$2.86 98.58$\pm$0.52 34.89$\pm$3.05 94.18$\pm$2.04 29.39$\pm$2.60 89.11$\pm$3.27
+w/o autoreg 4x 42.95$\pm$2.44 96.25$\pm$2.22 38.79$\pm$2.23 96.01$\pm$1.26 31.51$\pm$2.42 87.11$\pm$3.68
+8x 37.65$\pm$3.32 91.45$\pm$3.80 34.28$\pm$2.33 92.74$\pm$1.84 26.84$\pm$2.57 82.31$\pm$3.80
+w model agg 4x 46.65$\pm$2.86 99.58$\pm$0.14 38.52$\pm$3.62 95.50$\pm$2.32 33.14$\pm$2.52 92.80$\pm$2.75
 8x 40.80$\pm$3.12 97.96$\pm$0.82 32.94$\pm$3.52 91.83$\pm$3.59 28.77$\pm$2.23 87.85$\pm$3.43
+w synthetic 4x 46.56$\pm$2.84 99.56$\pm$0.15 38.80$\pm$3.55 96.31$\pm$1.72 33.44$\pm$2.72 93.37$\pm$2.37
+8x 40.41$\pm$3.02 98.43$\pm$0.75 33.99$\pm$3.38 93.70$\pm$2.57 28.47$\pm$2.57 87.63$\pm$3.63
 Across-site Site 1 (BraTS) Site 2 (fastMRI) Site 3 (MIDAS)
-reconstruction PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
-FedGAT 4x 36.71$\pm$3.01 95.02$\pm$1.96 40.28$\pm$2.65 96.39$\pm$1.81 43.52$\pm$2.83 98.06$\pm$1.04
-8x 32.02$\pm$2.95 91.53$\pm$3.07 35.52$\pm$2.84 93.68$\pm$2.55 37.98$\pm$2.97 95.65$\pm$1.94
-w model agg. 4x 35.83$\pm$3.12 94.15$\pm$2.54 39.89$\pm$2.70 96.19$\pm$1.95 42.59$\pm$3.26 97.54$\pm$1.64
+[…]
+FedGAT 4x 36.68$\pm$4.25 95.00$\pm$2.46 40.11$\pm$6.97 96.42$\pm$3.56 43.24$\pm$4.84 98.02$\pm$1.94
+8x 32.09$\pm$4.00 91.60$\pm$3.85 35.28$\pm$6.81 93.82$\pm$5.38 38.25$\pm$4.40 96.43$\pm$2.79
+w one-hot 4x 36.22$\pm$3.87 94.40$\pm$2.64 38.84$\pm$6.17 95.67$\pm$3.55 41.86$\pm$3.66 97.59$\pm$1.65
+prompt 8x 31.66$\pm$3.69 88.93$\pm$4.22 33.48$\pm$5.19 86.56$\pm$4.41 36.18$\pm$3.19 89.93$\pm$4.59
+w/o site 4x 36.34$\pm$4.33 94.50$\pm$2.75 39.99$\pm$6.85 96.18$\pm$3.86 42.85$\pm$4.85 97.67$\pm$2.33
+prompt 8x 31.56$\pm$3.92 90.49$\pm$4.14 35.21$\pm$6.49 93.38$\pm$5.36 37.48$\pm$4.61 95.39$\pm$3.28
+w/o scales 4x 34.16$\pm$3.88 91.25$\pm$3.97 38.14$\pm$6.92 94.76$\pm$5.29 42.11$\pm$4.55 97.50$\pm$2.28
+8x 30.33$\pm$3.91 87.18$\pm$5.31 34.22$\pm$6.37 92.58$\pm$5.79 36.76$\pm$4.31 95.28$\pm$3.22
+w/o autoreg 4x 35.15$\pm$4.32 91.56$\pm$5.23 37.23$\pm$6.21 91.68$\pm$5.49 40.87$\pm$3.13 96.13$\pm$1.81
+8x 30.56$\pm$4.46 87.52$\pm$6.01 32.25$\pm$6.17 86.88$\pm$5.95 35.97$\pm$3.33 92.10$\pm$3.06
+w model agg 4x 35.83$\pm$3.12 94.15$\pm$2.54 39.89$\pm$2.70 96.19$\pm$1.95 42.59$\pm$3.26 97.54$\pm$1.64
 8x 30.85$\pm$2.95 89.84$\pm$3.51 34.78$\pm$2.71 92.91$\pm$2.49 36.87$\pm$3.33 94.89$\pm$2.60
+w synthetic 4x 36.31$\pm$4.21 94.80$\pm$2.66 39.90$\pm$6.77 96.31$\pm$3.71 42.64$\pm$4.87 97.90$\pm$2.05
+8x 31.55$\pm$3.82 91.28$\pm$4.04 34.76$\pm$6.18 93.53$\pm$5.64 37.32$\pm$4.40 95.55$\pm$3.23
+5.3 Ablation Studies
+5.3.1 The GAT Prior
+FedGAT leverages synthetic MR images as a privacy-preserving means for cross-site knowledge transfer on local data distributions, and to enable collaborative training of generalizable reconstruction models in model-agnostic FL settings.
+To assess the utility of the proposed GAT prior in this context, we conducted a set of ablation studies comparing it against mainstream adversarial (GAN; [73]) and diffusion (DDPM; [84]) priors.
+Similar to the GAT prior, global GAN and DDPM priors were trained to capture the distribution of multi-site MRI data by conditioning them on a site prompt.
+First, we examined the fidelity of synthetic MR images produced by alternative priors in capturing the attributes of the underlying data distributions.
+Fidelity was measured via the FID metric between synthetic and actual MR images.
+FID values for GAT, GAN and DDPM priors are displayed in Fig.
+6, and sample synthetic MR images are shown in Fig.
+7.
+We find that GAT achieves the highest image fidelity, as manifested by its consistently lower FID scores compared to alternatives, and by its improved visual realism apparent in sample images.
+While GAN-generated images show some structural artifacts and suboptimal contrast and DDPM-generated images occasionally have residual noise, GAT achieves higher image quality with lower artifacts and noise.
+Second, we assessed whether the improvement in synthetic image fidelity achieved by the GAT prior translates onto benefits in reconstruction performance.
+For this purpose, performance of reconstruction models trained solely on synthetic images from the GAT, GAN or DDPM priors were compared, in reference to a benchmark model that was trained using actual multi-site MR images.
+Comparisons were performed in model-homogeneous settings (with MoDL at all sites) to isolate the effect of synthetic data quality and avoid confounding interactions between model architecture and data distribution across sites.
+As listed in Table 5, reconstruction models trained using the GAT prior yield significantly higher performance than those trained using GAN and DDPM priors (p$<$0.05), and they attain the most competitive performance against the benchmark model.
+These results corroborate that the superior image fidelity of the proposed GAT prior helps train more performant reconstruction models.
+As an FL technique, FedGAT communicates the parameters of the GAT prior instead of actual MRI data to mitigate potential risks in patient privacy.
+Yet, the prior might be susceptible to a level of information leakage across sites [24].
+Accordingly, we comparatively evaluated GAT, GAN and DDPM priors in terms of their susceptibility to information leakage.
+For this purpose, an intruder site was included in the original three-site FL setup (BraTS, fastMRI, MIDAS), with local dataset for the intruder site taken as random white-Gaussian noise images.
+Adopting a recent method to assess information leakage in FL, we examined the influence of the intruder site on the fidelity of synthetic images generated at remaining sites as an indication of information leakage [85].
+Table 6 lists fidelity of synthetic multi-site MR images generated based on GAT priors trained in the original three-site FL setup, and in the setup including the intruder site.
+We observe that GAN and DDPM priors experience a notable degree of loss in image fidelity, whereas the GAT prior shows relatively minute differences in fidelity.
+This result suggests that the GAT prior offers enhanced reliability against information leakage.
+5.3.2 Other Key Design Elements
+Next, we conducted another set of ablation studies to examine the contribution of key design elements in FedGAT to its performance.
+To assess the importance of utilizing a learnable site prompt via a gated MLP module, a ‘w one-hot prompt’ variant used an unlearned one-hot encoding of site index as the site prompt, and a ‘w/o site prompt’ variant entirely removed the site prompt from the input of the transformer module.
+To assess the importance of generating discrete token maps via a multi-scale autoregressive prediction, a ‘w/o scales’ variant ablated multi-scale generation of discrete token maps by setting $S=1$ (i.e., using a single spatial scale), and a ‘w/o autoreg’ variant ablated autoregressive prediction across scales by removing $f_{<S}$ from the input to the transformer module (i.e., predicting discrete token maps without knowledge of token maps at previous scales).
+To assess the importance of using hybrid local-synthetic datasets for cross-site knowledge transfer, a ‘w model agg’ variant was formed that simultaneously transferred knowledge and built reconstruction models via conventional model aggregation [37], and a ‘w synthetic’ variant was formed that built site-specific reconstruction models exclusively on synthetic datasets.
+The comparisons were again performed in model-homogeneous settings to avoid confounding interactions between model architecture and data distribution across sites.
+Performance metrics for variant models are listed in Table 7.
+FedGAT yields superior performance against the variant models in all cases (p$<$0.05).
+These results indicate that the individual design elements in FedGAT pertaining to the use of a gated-MLP driven site prompt, autoregressive and multi-scale generation of feature maps, and cross-site knowledge transfer via image synthesis all contribute significantly to its reconstruction performance.
+6 Discussion
+Here we introduced a model-agnostic FL technique for MRI reconstruction that supports model-heterogeneous settings by decoupling cross-site knowledge transfer from building of reconstruction models.
+In particular, FedGAT mediates knowledge transfer among sites by sharing a global GAT prior that uses autoregressive predictions across multiple spatial scales to generate synthetic MR images, and a site prompt for precise control over the synthetic image distribution.
+Previous imaging studies aiming to build generative priors have predominantly adopted adversarial approaches that can suffer from poor image quality due to training instabilities [19, 17, 86], or diffusion approaches that can suffer from extensive inference times and residual image noise [87, 88, 89, 90].
+In comparison, the autoregressive nature of GAT can offer improved image fidelity over adversarial and diffusion priors in FL settings as indicated by our results.
+To curate a hybrid multi-site MRI dataset, the shared GAT prior is employed at each site to efficiently generate high-fidelity synthetic MRI datasets from remaining sites.
+Afterwards, site-specific reconstruction models are locally trained on this hybrid dataset to improve generalization while enjoying freedom in architecture selection.
+Therefore, the proposed method holds great promise for advancing the practicality and scalability of FL in multi-institutional MRI studies.
+6.1 Site Specialization
 The distribution of MRI datasets collected in separate sites inevitably reflects the variations in imaging protocols or scanners across sites [91].
@@
 In turn, FedGAT can sustain a higher-degree of site specialization in reconstruction models than permitted by previous approaches.
-Devised to accurately capture an underlying distribution, generative priors synthesize MR images depicting random anatomical structures that do not correspond to the precise anatomy of any actual subject [75, 80].
+6.2 Computational Considerations
+FedGAT might offer a unique set of computational advantages for building MRI reconstruction models compared to existing FL frameworks.
+In conventional FL, sites must agree on a reconstruction model of common architecture, and each communication round involves training and aggregating local copies of this model.
+If a different architecture is later preferred, the entire FL training must be repeated, incurring substantial computational costs for each imaging site [30].
+FedGAT avoids this overhead by training a shared generative prior to model the distribution of multi-site MRI data independently of reconstruction.
+Once trained, the prior can be reused to synthesize data for locally building reconstruction models of various architectures, amortizing the cost of FL training across experiments on multiple architectures.
+In distillation-based FL methods, model heterogeneity is supported by maintaining separate site-specific and global reconstruction models, and aligning between these models through repeated distillation steps on the server and/or at individual sites [75].
+FedGAT bypasses this overhead by enabling cross-site knowledge transfer through synthetic image generation, eliminating the need for direct model alignment.
+While previous generative FL methods also adopt a decoupled training approach based on adversarial or diffusion priors [59], FedGAT offers a refined balance across sample fidelity and inference cost.
+Our results indicate that the GAT prior achieves higher image fidelity than both adversarial and diffusion priors, while requiring moderate inference time—faster than diffusion (10 autoregressive stages vs.
+1000 diffusion steps) and slower than adversarial methods (which require just one step).
+6.3 Fidelity of Synthetic Images
+Devised to capture an underlying distribution, generative priors synthesize MR images depicting random anatomical structures that do not correspond to the precise anatomy of any actual subject [87, 92].
 In FedGAT, such synthetic MR images from separate sites are employed in training of site-specific reconstruction models to improve generalization across sites.
-Our assessments based on FID scores indicate that the synthetic MR images generated by the GAT prior attain a level of quality and realism closely mimicking actual MR images.
+Our assessments based on FID scores indicate that the synthetic MR images generated by the GAT prior attain a level of quality and diversity closely mimicking actual MR images.
 Furthermore, reconstruction models trained on these synthetic MR images yield high performance when independently evaluated on actual MRI acquisitions in the test sets.
@@
 It remains important future work to validate the proposed method and its anatomical fidelity on broader patient cohorts.
+6.4 Patient Privacy
 FL frameworks help mitigate patient privacy risks by exchanging model weights rather than actual MRI data across sites.